### 11 Mar 2026 v1

# Surrogate-Assisted Targeted Learning for Delayed Outcomes under Administrative Censoring

Abstract
Delayed primary outcomes and administratively censored follow-up create a general semiparametric estimation problem: the target causal functional depends on an endpoint observed only for a shrinking subset of units at analysis time, while earlier surrogate measurements remain widely available. In such settings, inverse-probability-weighted estimators can become unstable as observation probabilities approach the positivity boundary, and complete-case model-based analyses can be highly sensitive to outcome-model specification.
We develop a surrogate-assisted targeted minimum loss estimator for this nested causal functional. Identification proceeds through a surrogate-bridge representation that integrates an observed-outcome regression over the conditional surrogate distribution, thereby avoiding inverse observation weights in the target parameter itself. We show that the estimator is asymptotically linear and doubly robust (in the sense that first-order bias vanishes when either nuisance component is consistently estimated), and we characterize two structural features of the problem: under surrogate-mediated missing at random, the censoring mechanism contributes no separate tangent-space component to the efficient influence function; and for nested bridge functionals, a one-step debiased machine-learning construction leaves a second-order cross-product remainder involving the conditional surrogate law. The proposed two-stage targeting step removes this term without requiring direct estimation of that law.
Simulation studies demonstrate stable finite-sample performance under substantial administrative censoring, and a design-calibrated analysis based on the Washington State EPT study illustrates the method in a realistic stepped-wedge cluster-randomized setting.
1 Introduction
Many modern studies measure the primary endpoint only after substantial delay, while shorter-term surrogate outcomes are available much earlier in follow-up. When analysis occurs before all primary outcomes mature, the resulting observed-data structure poses a distinctive semiparametric problem: the estimand depends on a delayed endpoint observed only for a subset of units whose observation probability may approach the positivity boundary, yet a surrogate remains broadly observed across the design space. In such settings, estimators that rely directly on inverse observation weights can become unstable as observation probabilities become small, whereas complete-case model-based analyses may inherit strong sensitivity to outcome-model specification.
This paper studies that problem through a surrogate-bridge formulation for causal effect estimation under administrative censoring. The key idea is to identify the target functional — the average treatment effect $\Psi(P_{0})=\mathbb{E}[Y(1)-Y(0)]$ — by integrating an observed-outcome regression over the conditional surrogate law, rather than by placing the inverse observation probability directly in the target parameter. This perspective leads naturally to a targeted learning estimator and exposes new structure in the corresponding semiparametric theory.
A stepped-wedge cluster randomized trial (SW-CRT) provides a particularly sharp motivating example. Because treatment rollout is tied to calendar time, late-crossing clusters may have little opportunity to realize the delayed endpoint before study closure. But the underlying statistical issue is broader than stepped-wedge designs: whenever an early surrogate is observed widely and a primary endpoint is administratively censored at analysis time, the analyst faces the same nested surrogate-bridge problem. We therefore use the stepped-wedge setting as a concrete and practically important instance of a more general semiparametric estimation problem.
Limitations of standard approaches.
Parametric mixed models (Hussey and Hughes, 2007; Hemming and Taljaard, 2020; Hughes et al., 2020) restrict attention to the complete-case sub-cohort $\{\Delta_{ijt}=1\}$ and require correct specification of both the secular time trend and the missing-data mechanism. IPCW estimators (Robins et al., 1994; Bang and Robins, 2005) place $g_{\Delta}^{-1}$ directly in the estimating equation; even moderate concentration of $g_{\Delta}$ near zero for late-crossing clusters produces severe variance inflation, and in the near-boundary regime these estimators may fail to admit regular asymptotic approximations.
Contributions.
This paper makes three methodological contributions.
First, we identify a class of causal functionals under surrogate-mediated administrative censoring through a nested bridge representation (Theorem 1) that avoids inverse observation weights in the target parameter. The surrogate-bridge G-computation formula identifies $\Psi(P_{0})$ via nested conditional expectations over the observed surrogate distribution, replacing the unstable $g_{\Delta}^{-1}$ with a support positivity condition on the complete-case outcome regression.
Second, we characterize the efficient-score structure of this problem (Section 3): under surrogate-mediated missing at random, the censoring mechanism contributes no separate tangent-space component to the efficient influence function (Lemma 1), so estimating $g_{\Delta}$ does not reduce the efficiency bound; and under clustered dependence, valid inference requires aggregation of influence contributions at the cluster level rather than averaging (Lemma 2).
Third, we show that for this nested functional, a standard one-step debiased machine-learning construction generally leaves a second-order cross-product remainder $R_{SY}$ involving the conditional surrogate law $f_{S}$ (Proposition 2); the proposed two-stage targeted minimum loss procedure removes that term without requiring direct estimation of $f_{S}$, achieving $\sqrt{J}$-consistency under a product-rate condition on the remaining nuisance estimators (Theorem 2).
Related statistical work.
Our work sits at the intersection of semiparametric missing-data theory, targeted learning, and modern debiased machine learning. Classical AIPW theory (Robins et al., 1994; Bickel et al., 1993) characterizes efficient estimation with missing outcomes via tangent space decomposition, but does not address the nested bridge structure considered here. Existing SW-CRT methodology (Hussey and Hughes, 2007; Hemming and Taljaard, 2020; Kasza et al., 2019; Hughes et al., 2020) largely remains centered on parametric mixed models and does not accommodate the present surrogate-assisted functional. The surrogate outcome literature (Gilbert et al., 2008; Frangakis and Rubin, 2002) addresses a complementary problem — evaluating treatment effects when the primary outcome is rare, not delayed — and their identification strategies do not extend to the near-boundary censoring regime. Luedtke et al. (2017) construct the closest prior estimator: a semiparametric surrogate-index estimator for i.i.d. data whose identification formula is structurally related to ours. However, their asymptotic linearity argument requires a uniform positivity lower bound on $g_{\Delta}$, which the SW-CRT’s calendar-driven censoring structure violates for late-crossing clusters; the present surrogate-bridge approach does not place $g_{\Delta}^{-1}$ in the target parameter and therefore does not require this condition. Cluster-level targeted learning (Balzer et al., 2016, 2021) establishes the importance of clusterwise influence aggregation for valid sandwich inference in group-randomized trials, but without the bridge-based identification and nested targeting problem studied here. The debiased machine learning framework (Chernozhukov et al., 2018) provides general product-rate guarantees via sample splitting, yet in the present setting a one-step construction leaves a nested cross-product remainder $R_{SY}$ involving the conditional surrogate law $f_{S}$ that cross-fitting does not eliminate. The contribution of this paper is to isolate that remainder explicitly (Proposition 2) and show how a second targeting step absorbs it without estimating $f_{S}$. [^1] See Kennedy (2022) for a comprehensive review of debiased machine learning. Proposition 2 gives the formal derivation.
Section 2 defines the observed data structure and establishes non-parametric identification via the surrogate-bridge representation. Section 3 develops the semiparametric theory: the EIC decomposition, the vanishing censoring-mechanism component (Lemma 1), the cluster-level summation rule (Lemma 2), and the nested cross-product remainder that makes two-stage targeting necessary (Proposition 2). Section 4 constructs the SA-TMLE as the direct solution to the theoretical constraints identified in Section 3. Section 5 establishes asymptotic linearity, double robustness, and the Berry–Esseen finite-sample coverage bound. Section 6 reports Monte Carlo experiments across three scenario blocks. Section 7 provides a brief design-calibrated illustration using the Washington State EPT trial.
Although our motivating application is a stepped-wedge cluster trial, the framework is not tied to that design. The same statistical structure arises whenever a delayed primary outcome is observed only for a subset of units at analysis time, an earlier surrogate is broadly available, and identification proceeds through a nested bridge representation. The paper is best understood not as a stepped-wedge method but as a general semiparametric framework for nested surrogate-bridge inference under administrative censoring, with stepped-wedge trials providing a particularly consequential use case.
Table 1: When to use each estimator: a practical comparison for stepped-wedge trials with delayed outcomes.
GLMM IPCW/AIPW SA-TMLE (proposed)
Target parameter Conditional treatment effect (model-based) Population ATE Population ATE
Key assumptions Correct outcome model specification (secular trend, random effects) Observation positivity: $g_{\Delta}^{-1}$ must be well-behaved Support positivity for $E[Y\mid S,A,W,t]$; surrogate-mediated MAR
Handles heavy censoring? No: complete-case analysis drops late-crossing clusters entirely Poorly: variance inflates as $g_{\Delta}\to 0$ Yes: bypasses $g_{\Delta}^{-1}$ via surrogate bridge
Requires surrogate? No No Yes
Double robustness? No Yes (but requires bounded $g_{\Delta}^{-1}$) Yes (asymptotically; finite-sample behavior may be asymmetric under nuisance misspecification)
Nonparametric nuisance? No (parametric outcome model) Yes (machine learning compatible) Yes (machine learning compatible)
Best suited for Low censoring, well-understood secular trends, no surrogate available Moderate censoring with good observation-positivity overlap Substantial administrative censoring with an available early surrogate
2 Data Structure and Identification
We first describe the observed-data structure and identification strategy in the stepped-wedge cluster-randomized setting, which serves as our running example of a broader class of surrogate-assisted problems with administratively censored delayed outcomes.
Consider a SW-CRT with $J$ clusters, $T$ calendar time steps, and $n_{j}$ individuals in cluster $j$ ($N=\sum_{j}n_{j}$ total). Each cluster crosses over once at a design-determined time $\tau_{j}\in\{2,\ldots,T\}$, so $A_{jt}=I(t\geq\tau_{j})$; crossover times are assigned uniformly at random and $A_{j1}=0$ for all $j$. The observed data tuple is
$O_{ijt}=\bigl(W_{ijt},\;A_{jt},\;S_{ijt},\;\Delta_{ijt},\;\Delta_{ijt}\cdot Y_{ijt}\bigr)\sim P_{0},$ (1)
where $W_{ijt}\in\mathbb{R}^{p}$ are baseline covariates; $S_{ijt}$ is the short-term surrogate; $\Delta_{ijt}\in\{0,1\}$ indicates whether the delayed primary outcome was observed by the administrative end date; and $Y_{ijt}$ is that outcome. The product $\Delta_{ijt}\cdot Y_{ijt}$ makes explicit that $Y$ is structurally missing for late-crossing clusters.
Let $\mathbf{U}=(U_{W},\,U_{A},\,b_{j},\,U_{S},\,U_{\Delta},\,U_{Y})$ denote the exogenous noise variables. The cluster-level random effect $b_{j}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\sigma^{2}_{b})$ is i.i.d. across clusters. All other components are individual-level noise terms assumed mutually independent across individuals and clusters.
The following system defines the endogenous variables in their causal time-ordering:
$W_{ijt}$ $=f_{W}(U_{W,ij}),$ (baseline covariates) (2)
$A_{jt}$ $=I(t\geq\tau_{j}),\quad\tau_{j}=f_{A}(U_{A,j}),$ (treatment; deterministic given $\tau_{j}$) (3)
$S_{ijt}$ $=f_{S}(W_{ijt},\,A_{jt},\,b_{j},\,t,\,U_{S,ijt}),$ (short-term surrogate) (4)
$\Delta_{ijt}$ $=f_{\Delta}(S_{ijt},\,A_{jt},\,W_{ijt},\,t,\,U_{\Delta,ijt}),$ (observation indicator) (5)
$Y_{ijt}$ $=f_{Y}(W_{ijt},\,A_{jt},\,S_{ijt},\,b_{j},\,t,\,U_{Y,ijt}).$ (delayed primary outcome) (6)
Three structural features deserve emphasis. First, $A_{jt}$ is deterministic given $\tau_{j}$; the treatment mechanism is known by design and does not require data-adaptive estimation (see Remark 4). Second, the cluster random effect $b_{j}$ enters both $f_{S}$ and $f_{Y}$, inducing the ICC among all observations within cluster $j$ (Remark 1). Third — and most critically — $Y$ does not appear in $f_{\Delta}$ and $U_{\Delta}$ does not appear in $f_{Y}$. This is the structural encoding of Assumption 3: conditional on $S$, the censoring indicator $\Delta$ is independent of the outcome $Y$. In DAG terms, there is no directed edge $Y_{ijt}\to\Delta_{ijt}$ (Figure 1).
[…]
Figure 1: Directed Acyclic Graph for the SW-CRT Structural Causal Model. Red circles: endogenous variables. Blue circles: exogenous and design variables. Violet arrows: cluster random effect $b_{j}$, the source of ICC. The critical structural feature is the absence of a directed edge $Y_{ijt}\!\to\!\Delta_{ijt}$, encoding Assumption 3: once $S_{ijt}$ is observed, the censoring probability depends on $S$ but not directly on the unobserved $Y$. The secular time trend $t$ acts as a common cause of all endogenous variables and must be adjusted for in estimation.
Remark 1 (Intra-cluster correlation).
The shared cluster random effect $b_{j}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\sigma^{2}_{b})$ enters both $f_{S}$ and $f_{Y}$, inducing ICC $\sigma^{2}_{b}/(\sigma^{2}_{b}+\sigma^{2}_{\varepsilon})$ among individuals within a cluster. The two methodological consequences are: (i) inference must use the cluster-robust sandwich estimator of Section 3; and (ii) Super Learner cross-validation must be performed at the cluster level (Supplement, Section S3). The TMLE estimator itself makes no parametric assumption about $b_{j}$; the Gaussian model governs the simulation DGP only.
2.1 Potential Outcomes and the Causal Target
For intervention value $a\in\{0,1\}$, let $S_{ijt}(a)$ and $Y_{ijt}(a)$ denote the potential outcomes obtained by the intervention $\mathrm{do}(A_{jt}=a)$ in the SCM. The causal Average Treatment Effect (ATE) is:
$\Psi(P_{0})\;=\;\mathbb{E}\bigl[Y_{ijt}(1)-Y_{ijt}(0)\bigr],$ (7)
where the expectation is over the marginal distribution of $(W_{ijt},b_{j},t)$ in the trial. The identification of $\Psi(P_{0})$ from $P_{0}$ is the subject of Section 2.2.
2.2 Non-Parametric Identification
The central identification insight is that the target causal effect can be represented through a surrogate bridge: one first learns the observed-outcome regression among units with realized primary outcomes and then integrates that regression over the treatment-specific surrogate distribution. This representation identifies the causal estimand without placing inverse observation weights in the target functional itself, which is precisely what makes it attractive in regimes where observation probabilities become small.
We establish the following assumptions to formalize this representation.
Assumption 1 (Consistency).
For any observed treatment assignment $A_{jt}=a$,  $S_{ijt}=S_{ijt}(a)$ and $Y_{ijt}=Y_{ijt}(a)$.
Assumption 2 (Sequential Randomization).
Conditional on the cluster history, treatment is independent of potential outcomes:
$\bigl(Y(0),Y(1),S(0),S(1)\bigr)\;\perp\;A_{jt}\;\big|\;W_{ijt},\,t.$ (8)
This holds by design: $A_{jt}=I(t\geq\tau_{j})$ and $\tau_{j}$ is randomized independently of individual outcomes.
Assumption 3 (Surrogate-Mediated Missing at Random).
The unobserved outcome provides no additional information about the censoring probability once the surrogate is known (Rubin, 1976):
$P(\Delta_{ijt}=1\mid Y_{ijt},S_{ijt},A_{jt},W_{ijt},t)\;=\;P(\Delta_{ijt}=1\mid S_{ijt},A_{jt},W_{ijt},t).$ (9)
This is structurally encoded in the SCM by the absence of the edge $Y_{ijt}\to\Delta_{ijt}$ (Figure 1, Equation (5)).
Remark 2 (Practical plausibility of Assumption 3).
Assumption 3 is most defensible when missingness is administrative (calendar-driven) rather than informative (health-driven). In the SW-CRT setting, the dominant censoring mechanism — the trial’s fixed end date — is independent of $Y_{ijt}$ conditional on $S_{ijt}$. A strong surrogate candidate satisfies three properties: (i) causal proximity: $S$ lies on or near the pathway $A\to S\to Y$; (ii) temporal availability: $S$ is observed for all individuals regardless of crossover time; and (iii) biological stationarity: $\mathbb{E}[Y\mid S,A,W,t]$ does not change systematically over calendar time. Examples include 3-month viral load predicting 12-month suppression in HIV trials, and discharge-process scores predicting 30-day readmission in quality-improvement SW-CRTs.
Assumption 4 (Support Positivity for the Observed Outcome Regression).
There exists a constant $c>0$ such that, for all $(s,a,w,t)$ in the support of $(S,A,W,t)$ relevant to the target parameter,
$P(\Delta_{ijt}=1\mid S_{ijt}=s,A_{jt}=a,W_{ijt}=w,t)\;\geq\;c.$ (10)
Remark 3 (Support positivity versus inverse-weighting regularity).
Assumption 4 is a support condition for identifying $E[Y\mid S,A,W,t]$ from the observed cohort $\{\Delta=1\}$, not an inverse-weighting regularity condition. IPCW-type estimators are sensitive to small values of $g_{\Delta}$ because $g_{\Delta}^{-1}$ appears directly in the influence function, so that even when Assumption 4 holds with a small constant $c$, the IPCW estimating equation becomes highly variable. The surrogate-bridge parameter, by contrast, does not contain $g_{\Delta}^{-1}$ in the target functional; it requires only that the complete-case outcome regression $E[Y\mid S,A,W,t,\Delta=1]$ be well-defined on the support over which the surrogate-bridge formula integrates.
Assumption 5 (Design-Based Treatment Support).
The marginal treatment probability, averaged over the trial design space, is non-degenerate:
$0\;<\;\mathbb{E}_{t}\bigl[g_{A}(1\mid W,t)\bigr]\;<\;1.$ (11)
Remark 4 (Why Treatment Positivity Holds Marginally Despite Pointwise Violations).
At any fixed time $t$ and cluster $j$, the treatment assignment is deterministic: $A_{jt}=I(t\geq\tau_{j})$, so pointwise positivity $0<P(A_{jt}=1\mid W,t)<1$ fails everywhere. A naive reading would therefore conclude that treatment positivity is structurally violated by the SW-CRT design.
This is resolved by observing that the identification formula (1) marginalizes over the trial’s design space, averaging jointly over both $W$ and $t$. The design-based propensity is known by construction from the randomization schedule. Under a uniform randomization of crossover times:
$g_{A}(1\mid W,t)\;=\;P(\tau_{j}\leq t)\;=\;\frac{t-1}{T-1}.$ (12)
Averaging over the interior steps $t\in\{2,\ldots,T-1\}$ yields $\mathbb{E}_{t}[g_{A}(1\mid W,t)]=1/2$, so Assumption 5 holds. In practice, boundary steps $t=1$ and $t=T$ should be excluded, as $g_{A}$ equals 0 and 1 respectively at these endpoints.
A further consequence is that $g_{A}$ need not be estimated from the data. In Section 4, we plug in the design-based propensity directly, eliminating one source of nuisance estimation error from the second-order remainder $R_{2}$ (Supplement, Section S2).
Theorem 1 (Identification via a Surrogate Bridge under Support Positivity).
Under Assumptions 15, the causal ATE is identified by the longitudinal G-computation formula:
$\Psi(P_{0})\;=\;\mathbb{E}_{W,t}\Bigl($ $\mathbb{E}_{S\mid A=1,W,t}\bigl[\mathbb{E}[Y\mid S,A=1,W,t,\Delta=1]\bigr]$
$-\;\mathbb{E}_{S\mid A=0,W,t}\bigl[\mathbb{E}[Y\mid S,A=0,W,t,\Delta=1]\bigr]\Bigr).$ (13)
Proof.
By Assumptions 1 (Consistency) and 2 (Sequential Randomization), the potential outcome mean satisfies:
$\mathbb{E}[Y(a)\mid S(a),W,t]\;=\;\mathbb{E}[Y\mid S,A=a,W,t].$ (14)
By Assumption 3 (Surrogate-Mediated MAR) and Assumption 4 (Support Positivity), we may condition on the observed cohort ($\Delta=1$):
$\mathbb{E}[Y\mid S,A=a,W,t]\;=\;\mathbb{E}[Y\mid S,A=a,W,t,\Delta=1]\;\equiv\;\bar{Q}_{Y}(S,a,W,t).$ (15)
Integrating $\bar{Q}_{Y}$ over the conditional surrogate distribution:
$\mathbb{E}[Y(a)\mid W=w,t]\;=\;\int\bar{Q}_{Y}(s,a,w,t)\,dP(s\mid a,w,t)\;\equiv\;\bar{Q}_{\mathrm{int}}(a,w,t).$ (16)
Marginalizing over the trial design space yields (1).∎∎
3 Semiparametric Theory
We now study the surrogate-bridge estimand as a semiparametric functional. This perspective is central for two reasons. First, it reveals the efficient-score geometry induced by surrogate-mediated censoring and clustered dependence: the censoring mechanism contributes no separate tangent-space component, and the cluster structure requires summation rather than averaging of influence contributions. Second, it clarifies why a standard one-step debiased estimator is not sufficient for the nested bridge problem: the second-order remainder contains a cross-product term involving the conditional surrogate law that is not removed by ordinary cross-fitting. The resulting analysis motivates the specific two-stage targeting construction developed in Section 4.
3.1 EIC Decomposition
The individual-level EIC components, whose derivation is provided in Supplement, Section S2, are as follows. For treatment $a\in\{0,1\}$:
$D_{Y,a}(O_{ijt})$ $=\frac{I(A=a)\,\Delta}{g_{A}(a\mid W,t)\,g_{\Delta}(1\mid S,A,W,t)}\bigl(Y-\bar{Q}_{Y}(S,\,a,\,W,t)\bigr),$ (17)
$D_{S,a}(O_{ijt})$ $=\frac{I(A=a)}{g_{A}(a\mid W,t)}\bigl(\bar{Q}_{Y}(S,\,a,\,W,t)\;-\;\bar{Q}_{\mathrm{int}}(a,\,W,t)\bigr),$ (18)
$D_{W,a}(O_{ijt})$ $=\bar{Q}_{\mathrm{int}}(a,W,t)-\hat{\Psi}_{a},$ (19)
where $\hat{\Psi}_{a}=N^{-1}\sum_{j,i}\bar{Q}^{\,*}_{\mathrm{int}}(a,W_{ijt},t_{ij})$.
In Equation (18), both nuisance functions are evaluated at the fixed intervention value $a$ rather than the observed random variable $A$. Under the indicator $I(A=a)$, the two are numerically equal on the support of the data, but the fixed-argument form is essential: it makes explicit that $\bar{Q}_{Y}(S,a,W,t)$ and $\bar{Q}_{\mathrm{int}}(a,W,t)$ are counterfactual predictions under $\mathrm{do}(A=a)$, consistent with the identification formula (1). Writing $A$ in the argument of $D_{S,a}$ would obscure this causal interpretation and introduce an inconsistency with the identification strategy.
The total individual EIC is:
$D^{*}(O_{ijt})\;=\;\bigl(D_{Y,1}+D_{S,1}+D_{W,1}\bigr)\;-\;\bigl(D_{Y,0}+D_{S,0}+D_{W,0}\bigr).$ (20)
The first structural result shows that surrogate-mediated censoring changes the identification strategy but not the efficient-score geometry in the usual missingness direction: once the surrogate is conditioned upon, the censoring mechanism contributes no separate tangent-space component.
Lemma 1 (Vanishing $\mathcal{T}_{\Delta}$ Component).
Under Assumption 3 (Surrogate-Mediated MAR), the pathwise derivative of $\Psi_{a}$ in the $\mathcal{T}_{\Delta}$ direction is identically zero: the EIC has no censoring-mechanism component.
Proof.
Consider a parametric submodel perturbing only $g_{\Delta}$ along score $h_{\Delta}\in\mathcal{T}_{\Delta}$. The identification formula (Equation (1)) reaches $g_{\Delta}$ only through the restriction of $\mathbb{E}[Y\mid S,A{=}a,W,t]$ to $\{\Delta=1\}$. By Assumption 3, $P(\Delta=1\mid Y,S,A,W,t)=P(\Delta=1\mid S,A,W,t)$, so perturbing $g_{\Delta}$ while holding $\bar{Q}_{Y}(S,a,W,t)$ and $P(S\mid a,W,t)$ fixed leaves $\bar{Q}_{\mathrm{int}}(a,W,t)$ unchanged. Consequently,
$\frac{d}{d\epsilon}\Psi_{a}(P_{\epsilon})\big|_{\epsilon=0}=\mathbb{E}_{P}\!\bigl[\bigl(\bar{Q}_{\mathrm{int}}(a,W,t)-\Psi_{a}\bigr)\,h_{\Delta}(O)\bigr]=0,$ (21)
where the final equality holds because $\bar{Q}_{\mathrm{int}}(a,W,t)-\Psi_{a}$ is $P$-mean-zero and orthogonal to any score supported only on variations in the censoring mechanism. Without Assumption 3, the censoring direction contributes a non-zero term involving the unidentified quantity $\mathbb{E}[Y\mid S,A,W,t,\Delta=0]$, rendering the EIC non-estimable from observed data. ∎∎
The second structural result concerns the correct level of influence aggregation. Because the estimand is defined through cluster-randomized data with arbitrary within-cluster dependence, the efficient contribution of a cluster is obtained by summation rather than averaging of individual terms; this is essential for valid semiparametric variance estimation under unequal cluster sizes.
Lemma 2 (Cluster-Level Summation Rule).
Let $O_{j}=\{O_{ijt}:i=1,\ldots,n_{j},\;t=1,\ldots,T\}$ denote the full data vector for cluster $j$, and let $\Psi$ be expressed as a functional of the cluster-level distribution $P_{j}$. The Riesz representer of the pathwise derivative in $L^{2}(P_{j})$ — equivalently, the cluster-level EIC — is the sum (not average) of individual EICs:
$\mathrm{EIC}_{j}=\sum_{i=1}^{n_{j}}\sum_{t=1}^{T}D^{*}(O_{ijt}).$ (22)
Averaging by $n_{j}T$ would rescale the influence curve, rendering the sandwich variance estimator inconsistent under unbalanced cluster sizes and non-negligible ICC.
Proof.
Collecting Proposition 3.1 for $a=1$ minus $a=0$:
$D^{*}(O_{ijt})\;=\;\bigl(D_{Y,1}+D_{S,1}+D_{W,1}\bigr)-\bigl(D_{Y,0}+D_{S,0}+D_{W,0}\bigr).$ (23)
Because clusters are the i.i.d. units of randomization, the parameter $\Psi$ is a functional of the cluster-level distribution $P_{j}$ of $O_{j}=\{O_{ijt}:i=1,\ldots,n_{j},\,t=1,\ldots,T\}$:
$\Psi=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{n_{j}}\bigl[\bar{Q}_{\mathrm{int}}(1,W_{ij},t_{ij})-\bar{Q}_{\mathrm{int}}(0,W_{ij},t_{ij})\bigr].$ (24)
Applying the pathwise derivative to a submodel $P_{j,\epsilon}$ perturbing only cluster $j$ along score $h_{j}(O_{j})$, the chain rule over the additive structure gives
$\frac{d}{d\epsilon}\Psi(P_{j,\epsilon})\big|_{\epsilon=0}=\mathbb{E}_{P_{j}}\!\!\left[\left(\sum_{i=1}^{n_{j}}\sum_{t=1}^{T}D^{*}(O_{ijt})\right)h_{j}(O_{j})\right].$ (25)
The Riesz representer in $L^{2}(P_{j})$ is therefore the sum (not average) of individual EICs:
$\mathrm{EIC}_{j}\;=\;\sum_{i=1}^{n_{j}}\sum_{t=1}^{T}D^{*}(O_{ijt}).$ (26)
Summation absorbs both the intra-cluster correlation (ICC) across individuals at a given step and the serial correlation across steps induced by the secular trend; averaging by $n_{j}$ would rescale the influence curve and produce inconsistent variance estimates under unbalanced cluster sizes. ∎
3.2 The Nested Cross-Product Remainder and the Necessity of Two-Stage Targeting
The previous two lemmas establish the efficient-score structure of the surrogate-bridge functional — the core semiparametric object of this paper — but they do not yet determine the appropriate estimator. A natural question is whether a standard one-step debiased machine-learning construction suffices. The next result shows that, for nested bridge functionals, the answer is generally no. Even after cross-fitting removes first-order empirical-process terms, a second-order cross-product remainder $R_{SY}$ remains because the bridge functional depends jointly on the outcome regression and the conditional surrogate law $f_{S}$. This remainder is not a fluctuation term and therefore is not eliminated by ordinary sample splitting. This is the semiparametric obstruction that the two-stage targeting step of Section 4 is specifically designed to resolve.
Proposition 1 (Nested Cross-Product Term in the DML Remainder).
Let $f_{S}(\cdot\mid a,W,t)\equiv P_{0}(S\in\cdot\mid A=a,W,t)$ denote the conditional density of the surrogate, and let $\hat{f}_{S}$ be any estimator of $f_{S}$ independent of the SA-TMLE nuisance estimators. Define the plug-in integral estimator $\hat{\bar{Q}}^{\,\mathrm{DML}}_{\mathrm{int}}(a,W,t)\equiv\int\hat{\bar{Q}}_{Y}(s,a,W,t)\,\hat{f}_{S}(s\mid a,W,t)\,d\mu(s)$, which constructs $\bar{Q}_{\mathrm{int}}$ as an integral of $\hat{\bar{Q}}_{Y}$ against $\hat{f}_{S}$.
For the DML one-step estimator $\hat{\Psi}^{\mathrm{DML}}=\hat{\Psi}^{\mathrm{DML}}_{\mathrm{plug\text{-}in}}+N^{-1}\sum_{j,i,t}D^{*}(O_{ijt};\hat{\bar{Q}}_{Y},\hat{\bar{Q}}^{\,\mathrm{DML}}_{\mathrm{int}},\hat{g}_{A},\hat{g}_{\Delta})$ based on this construction, the von Mises expansion yields:
$\hat{\Psi}^{\mathrm{DML}}-\Psi(P_{0})\;=\;\frac{1}{J}\sum_{j=1}^{J}\mathrm{EIC}_{j}\;+\;R_{2}^{\mathrm{TMLE}}(\hat{P},P_{0})\;+\;R_{SY}(\hat{P},P_{0})\;+\;o_{P}(J^{-1/2}),$ (27)
where $R_{2}^{\mathrm{TMLE}}$ is the TMLE remainder (Supplement, Section S2.4) and the nested cross-product remainder is
$R_{SY}(\hat{P},P_{0})\;=\;\sum_{a\in\{0,1\}}(-1)^{1-a}\iint\bigl(\hat{\bar{Q}}_{Y}-\bar{Q}^{0}_{Y}\bigr)(s,a,w,t)\,\bigl(\hat{f}_{S}-f^{0}_{S}\bigr)(s\mid a,w,t)\;d\mu(s)\,dP_{0}(w,t).$ (28)
For $R_{SY}=o_{P}(J^{-1/2})$ the DML estimator requires the additional product-rate condition
$\sum_{a\in\{0,1\}}\bigl\|\hat{\bar{Q}}_{Y}(\cdot,a,\cdot,\cdot)-\bar{Q}^{0}_{Y}(\cdot,a,\cdot,\cdot)\bigr\|_{P_{0}}\cdot\bigl\|\hat{f}_{S}(\cdot\mid a,\cdot,\cdot)-f^{0}_{S}(\cdot\mid a,\cdot,\cdot)\bigr\|_{P_{0}}\;=\;o_{P}(J^{-1/2}),$ (29)
which involves estimating the conditional density $f_{S}$ at rate $o_{P}(J^{-1/4})$ in $L^{2}(P_{0})$.
The SA-TMLE eliminates $R_{SY}$ without estimating $f_{S}$: the second fluctuation step (Section 4.2) enforces $N^{-1}\sum_{j,i,t}D_{S,a}(O_{ijt})=0$, which constrains $\hat{\bar{Q}}^{\,*}_{\mathrm{int}}$ to the propensity-weighted empirical conditional mean of $\hat{\bar{Q}}^{\,*}_{Y}$, thereby absorbing $R_{SY}$ into the efficient score without requiring knowledge of $f_{S}$. The proof is given in Supplement, Section S2.6.
Remark 5 (Why DML cross-fitting does not eliminate $R_{SY}$).
[…]
4 Surrogate-Assisted TMLE Construction
Proposition 2 identifies the obstacle to a one-step debiased estimator: the nested bridge functional generates a second-order remainder $R_{SY}$ involving the conditional surrogate law $f_{S}$ that cross-fitting does not eliminate. The estimator developed here resolves that obstacle directly. Rather than constructing the bridge by separately estimating the surrogate density, the proposed two-stage targeted minimum loss procedure absorbs $R_{SY}$ into the efficient score, yielding an estimator that preserves doubly robust asymptotic linearity without direct estimation of $f_{S}$.
4.1 Stage 1: Initial Estimation via Super Learner
We seek to estimate the population functional $\Psi(P_{0})$ defined in (7). We first construct initial estimates of the four nuisance functions via a Super Learner ensemble (Supplement, Section S3):
$\bar{Q}_{Y}(S,A,W,t)$ $=\mathbb{E}[Y\mid S,A,W,t,\Delta=1],$ (30)
$\bar{Q}_{\mathrm{int}}(A,W,t)$ $=\mathbb{E}\bigl[\bar{Q}_{Y}(S,A,W,t)\mid A,W,t\bigr],$ (31)
$g_{A}(1\mid W,t)$ $=P(A=1\mid W,t),$ (32)
$g_{\Delta}(1\mid S,A,W,t)$ $=P(\Delta=1\mid S,A,W,t).$ (33)
Remark 6 (Observed-Data Regression vs. Counterfactual Evaluation of $\bar{Q}_{\mathrm{int}}$).
Equation (31) defines $\bar{Q}_{\mathrm{int}}(A,W,t)$ as an observed-data regression of $\bar{Q}_{Y}$ on the observed treatment $A$. During estimation, this regression is fit on the full data using the observed $A\in\{0,1\}$. Counterfactual predictions are then obtained by evaluating the fitted model at the fixed intervention values $A=1$ and $A=0$ separately, yielding $\hat{\bar{Q}}_{\mathrm{int}}(1,W,t)$ and $\hat{\bar{Q}}_{\mathrm{int}}(0,W,t)$. These counterfactual evaluations enter the ATE estimator (37). The regression is fit on the observed scale and evaluated on the interventional scale; conflating these two steps would produce a biased plug-in estimator, which is precisely what the TMLE fluctuation step in Section 4.2 corrects.
As established in Remark 4, the treatment propensity $g_{A}(1\mid W,t)=(t-1)/(T-1)$ is known from the randomization schedule and is plugged in directly. Only $g_{\Delta}$ requires data-adaptive estimation.
4.2 Stage 2: The Nested Fluctuation Step
Clever covariate for the surrogate integration model.
$H_{S}(O_{ijt})\;=\;\frac{1}{\hat{g}_{A}(A_{\mathrm{obs}}\mid W,t)},$ (34)
with the second fluctuation update fit on the full data:
$\operatorname{logit}\!\bigl(\bar{Q}^{\,*}_{\mathrm{int}}(\varepsilon_{S})\bigr)\;=\;\operatorname{logit}\!\bigl(\hat{\bar{Q}}_{\mathrm{int}}(A_{\mathrm{obs}},W,t)\bigr)\;+\;\varepsilon_{S}\,H_{S}(O_{ijt}).$ (35)
Convergence criterion.
The fluctuation step solves the efficient score equation. We declare convergence when the empirical mean of the cluster-level EIC (Section 3) falls below the tolerance:
$\left|\frac{1}{J}\sum_{j=1}^{J}\mathrm{EIC}_{j}\right|\;\leq\;\frac{1}{\sqrt{N}\,\log N}.$ (36)
In practice a single one-step fluctuation is sufficient when the Super Learner initial estimates are well-specified; we iterate only if (36) is not met after the first step.
Point estimator.
The final TMLE point estimate is the empirical mean of the doubly-targeted counterfactual predictions:
$\hat{\Psi}_{\mathrm{TMLE}}\;=\;\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{n_{j}}\Bigl[\bar{Q}^{\,*}_{\mathrm{int}}(1,W_{ijt},t_{ij})\;-\;\bar{Q}^{\,*}_{\mathrm{int}}(0,W_{ijt},t_{ij})\Bigr].$ (37)
5 Asymptotic Theory
5.1 Cluster-Level Aggregation and Inference
We aggregate the individual influence curves to the cluster level by summing over all individuals and all time steps within cluster $j$:
$\mathrm{EIC}_{j}\;=\;\sum_{i=1}^{n_{j}}\sum_{t=1}^{T}D^{*}(O_{ijt}).$ (38)
The double summation over both $i$ and $t$ is essential. Because the cluster is the independent unit of the data-generating process, $\mathrm{EIC}_{j}$ must absorb all within-cluster dependence: both the ICC across individuals at a given time step and the serial correlation across time periods induced by the secular trend. The sandwich variance estimator below is then valid without any additional autocorrelation correction, provided $J$ is sufficient for the cluster-level CLT. The formal projection argument establishing that summation (not averaging) is the correct operation is provided in Supplement, Section S2.
The cluster-robust sandwich variance estimator is:
$\widehat{\mathrm{Var}}\!\left(\hat{\Psi}_{\mathrm{TMLE}}\right)\;=\;\frac{1}{J}\cdot\frac{1}{J-1}\sum_{j=1}^{J}\left(\mathrm{EIC}_{j}-\overline{\mathrm{EIC}}\right)^{2},$ (39)
and a Wald confidence interval at level $1-\alpha$ is
$\mathrm{CI}^{z}_{1-\alpha}\;=\;\hat{\Psi}_{\mathrm{TMLE}}\;\pm\;z_{\alpha/2}\,\sqrt{\widehat{\mathrm{Var}}\!\left(\hat{\Psi}_{\mathrm{TMLE}}\right)},$ (40)
where $z_{\alpha/2}$ is the upper $\alpha/2$ quantile of $\mathcal{N}(0,1)$. For the small-$J$ regime, Theorem S1 (Supplement, Section S1) establishes that replacing $z_{\alpha/2}$ by the $t_{J-1}$ quantile strictly reduces the undercoverage risk; we denote this interval
$\mathrm{CI}^{t}_{1-\alpha}(J)\;=\;\hat{\Psi}_{\mathrm{TMLE}}\;\pm\;t_{J-1,\,\alpha/2}\,\sqrt{\widehat{\mathrm{Var}}\!\left(\hat{\Psi}_{\mathrm{TMLE}}\right)},$ (41)
where $t_{J-1,\alpha/2}$ is the upper $\alpha/2$ quantile of Student’s $t$-distribution with J1J-1 degrees of freedom.
5.2 Asymptotic Normality
We now show that the proposed estimator solves the semiparametric estimation problem induced by the nested bridge functional: after the two-stage targeting step, the estimator is asymptotically linear with influence function equal to the cluster-level efficient score, and the remaining second-order term obeys the usual product-rate control. Asymptotic linearity implies that the cluster-robust sandwich variance estimator is consistent for the true sampling variance.
Theorem 2 (Asymptotic Linearity and Normality of the Surrogate-Assisted TMLE).
Suppose Assumptions 15 hold and the following regularity conditions are satisfied:
  1. (C1)
    (Cluster independence) The cluster-level data vectors $O_{j}=\{O_{ijt}:i=1,\ldots,n_{j},\;t=1,\ldots,T\}$ are independent across $j=1,\ldots,J$, with $n_{j}$ bounded above by a constant $\bar{n}<\infty$.
  2. (C2)
    (Bounded outcomes) $Y$, $S$, and all nuisance functions are uniformly bounded almost surely under $P_{0}$.
  3. (C3)
    (Product-rate condition) The nuisance estimators satisfy
    $\bigl\|\hat{\bar{Q}_{Y}}-\bar{Q}^{0}_{Y}\bigr\|_{P_{0}}\cdot\bigl\|\hat{g_{\Delta}}-g^{0}_{\Delta}\bigr\|_{P_{0}}\;+\;\bigl\|\hat{\bar{Q}_{\mathrm{int}}}-\bar{Q}_{\mathrm{int}}^{0}\bigr\|_{P_{0}}\cdot\bigl\|\hat{g_{A}}-g^{0}_{A}\bigr\|_{P_{0}}\;=\;o_{P}(J^{-1/2}),$ (42)
    where $\|\cdot\|_{P_{0}}$ denotes the $L^{2}(P_{0})$ norm. This condition is satisfied whenever each nuisance estimator converges at rate $o_{P}(J^{-1/4})$ in $L^{2}(P_{0})$.
Then the TMLE is asymptotically linear:
$\hat{\Psi}_{\mathrm{TMLE}}-\Psi(P_{0})\;=\;\frac{1}{J}\sum_{j=1}^{J}\mathrm{EIC}_{j}\;+\;o_{P}(J^{-1/2}),$ (43)
and, as $J\to\infty$:
$\sqrt{J}\,\bigl(\hat{\Psi}_{\mathrm{TMLE}}-\Psi(P_{0})\bigr)\;\to_{d}\;\mathcal{N}\!\bigl(0,\;\sigma^{2}\bigr),\qquad\sigma^{2}\;=\;\mathrm{Var}(\mathrm{EIC}_{j}).$ (44)
The cluster-robust sandwich estimator (39) is consistent for $\sigma^{2}/J$, so the 95% confidence interval (40) achieves asymptotically exact coverage.
Proof.
Asymptotic linearity (43) follows from the von Mises expansion of $\hat{\Psi}_{\mathrm{TMLE}}$ around $P_{0}$. The TMLE fluctuation step is constructed precisely to solve the efficient score equation $J^{-1}\sum_{j}\mathrm{EIC}_{j}=0$ (up to the convergence tolerance (36)), which eliminates the first-order bias term. The residual is the second-order remainder $R_{2}(\hat{P},P_{0})$ characterized in Supplement, Section S2.4. Under condition (C3), $R_{2}=o_{P}(J^{-1/2})$; this follows from the bilinear form of $R_{2}$ established in Supplement, Section S2.4 and the Cauchy–Schwarz inequality applied to each product of $L^{2}(P_{0})$ errors. When the CV-TMLE of Supplement, Section S3 is used, condition (C3) is achieved without requiring a Donsker condition on the nuisance function classes, because sample splitting eliminates the empirical process terms that would otherwise arise (Zheng and van der Laan, 2011). Asymptotic normality (44) and consistency of the sandwich variance estimator then follow from the Lindeberg–Feller CLT applied to the i.i.d. cluster scores $\{\mathrm{EIC}_{j}\}_{j=1}^{J}$; the complete argument is given in Supplement, Section S2.4.∎∎
Remark 7 (Double robustness of the asymptotic result).
The product-rate condition (C3) is doubly robust in the following sense (Bang and Robins, 2005; Tsiatis, 2006): it is satisfied if either (i) the outcome models $(\bar{Q}_{Y},\bar{Q}_{\mathrm{int}})$ are consistently estimated at any rate, or (ii) the propensity mechanisms $(g_{A},g_{\Delta})$ are consistently estimated at any rate. In particular, since $g_{A}$ is known by design in the SW-CRT (Remark 4), the condition reduces to requiring only that $\hat{\bar{Q}_{Y}}$ and $\hat{\bar{Q}_{\mathrm{int}}}$ converge at rate $o_{P}(J^{-1/4})$, a strictly weaker requirement than pointwise consistency.
Proposition 2 (Nested Cross-Product Term in the DML Remainder).
Let $f_{S}(\cdot\mid a,W,t)\equiv P_{0}(S\in\cdot\mid A=a,W,t)$ denote the conditional density of the surrogate, and let $\hat{f}_{S}$ be any estimator of $f_{S}$ independent of the SA-TMLE nuisance estimators. Define the plug-in integral estimator $\hat{\bar{Q}}^{\,\mathrm{DML}}_{\mathrm{int}}(a,W,t)\equiv\int\hat{\bar{Q}}_{Y}(s,a,W,t)\,\hat{f}_{S}(s\mid a,W,t)\,d\mu(s)$, which constructs $\bar{Q}_{\mathrm{int}}$ as an integral of $\hat{\bar{Q}}_{Y}$ against $\hat{f}_{S}$.
For the DML one-step estimator $\hat{\Psi}^{\mathrm{DML}}=\hat{\Psi}^{\mathrm{DML}}_{\mathrm{plug\text{-}in}}+N^{-1}\sum_{j,i,t}D^{*}(O_{ijt};\hat{\bar{Q}}_{Y},\hat{\bar{Q}}^{\,\mathrm{DML}}_{\mathrm{int}},\hat{g}_{A},\hat{g}_{\Delta})$ based on this construction, the von Mises expansion yields:
$\hat{\Psi}^{\mathrm{DML}}-\Psi(P_{0})\;=\;\frac{1}{J}\sum_{j=1}^{J}\mathrm{EIC}_{j}\;+\;R_{2}^{\mathrm{TMLE}}(\hat{P},P_{0})\;+\;R_{SY}(\hat{P},P_{0})\;+\;o_{P}(J^{-1/2}),$ (45)
where $R_{2}^{\mathrm{TMLE}}$ is the TMLE remainder (Supplement, Section S2.4) and the nested cross-product remainder is
$R_{SY}(\hat{P},P_{0})\;=\;\sum_{a\in\{0,1\}}(-1)^{1-a}\iint\bigl(\hat{\bar{Q}}_{Y}-\bar{Q}^{0}_{Y}\bigr)(s,a,w,t)\,\bigl(\hat{f}_{S}-f^{0}_{S}\bigr)(s\mid a,w,t)\;d\mu(s)\,dP_{0}(w,t).$ (46)
For $R_{SY}=o_{P}(J^{-1/2})$ the DML estimator requires the additional product-rate condition
$\sum_{a\in\{0,1\}}\bigl\|\hat{\bar{Q}}_{Y}(\cdot,a,\cdot,\cdot)-\bar{Q}^{0}_{Y}(\cdot,a,\cdot,\cdot)\bigr\|_{P_{0}}\cdot\bigl\|\hat{f}_{S}(\cdot\mid a,\cdot,\cdot)-f^{0}_{S}(\cdot\mid a,\cdot,\cdot)\bigr\|_{P_{0}}\;=\;o_{P}(J^{-1/2}),$ (47)
which involves estimating the conditional density $f_{S}$ at rate $o_{P}(J^{-1/4})$ in $L^{2}(P_{0})$.
The SA-TMLE eliminates $R_{SY}$ without estimating $f_{S}$: the second fluctuation step (Section 4.2) enforces $N^{-1}\sum_{j,i,t}D_{S,a}(O_{ijt})=0$, which constrains $\hat{\bar{Q}}^{\,*}_{\mathrm{int}}$ to the propensity-weighted empirical conditional mean of $\hat{\bar{Q}}^{\,*}_{Y}$, thereby absorbing $R_{SY}$ into the efficient score without requiring knowledge of $f_{S}$. The proof is given in Supplement, Section S2.6.
Remark 8 (Why DML cross-fitting does not eliminate $R_{SY}$).
[…]
5.3 Non-asymptotic Coverage Bound
Beyond asymptotic validity, the bridge-based estimator admits an explicit non-asymptotic coverage analysis at the cluster level. The resulting Berry–Esseen bound (Theorem S1, Supplementary Material) quantifies how bounded cluster scores, variance concentration, and nuisance-estimation remainder jointly determine finite-sample undercoverage, thereby linking the semiparametric theory directly to the small-$J$ regimes encountered in practice. Corollary S1 gives the minimum cluster count $J^{*}$ for a target coverage level; for the simulation DGP of Section 6, $J^{*}\approx 27$ for 1% maximum undercoverage. Corollary S2 (Supplement, Section S2.7) extends this to characterize how nuisance convergence rates govern the sandwich variance gap, showing that the constant coverage shortfall observed in Block I simulations is the finite-sample signature of estimation near the product-rate boundary.
6 Monte Carlo Simulation Study
The simulation study is designed to examine the main qualitative implications of the theory: relative stability of the surrogate-bridge estimator as observation probabilities become small (Blocks I and III), the finite-sample behavior of the asymptotic double-robustness theory under deliberate nuisance misspecification (Block II), and the relationship between coverage and the remainder-variance analysis of Section 5.3. The three scenario blocks are organized around increasing administrative censoring severity, misspecified nuisance models, and varying cluster counts. Throughout, “double robustness” refers to the asymptotic identification of the first-order remainder; finite-sample performance can still depend strongly on which nuisance component is misspecified and on the magnitude of the residual second-order error.
6.1 Data-Generating Process
Trial design.
We simulate a standard SW-CRT with $J$ clusters, $T=7$ calendar time steps, and $n_{j}=40$ individuals per cluster per step, giving a total sample size of $N=J\times T\times n_{j}$. Crossover times $\tau_{j}$ are drawn uniformly without replacement from $\{2,3,4,5,6,7\}$, so that exactly J/(T1)J/(T-1) clusters (rounded to the nearest integer) cross over at each step. All clusters begin in the control condition: $A_{jt}=\mathbf{1}(t\geq\tau_{j})$. Administrative censoring of the delayed outcome is structurally imposed: an individual in cluster $j$ at time step $t$ can contribute an observed delayed outcome only if $\tau_{j}+t_{\mathrm{lag}}\leq T+t_{\mathrm{grace}}$, where $t_{\mathrm{lag}}=2$ steps is the outcome delay and $t_{\mathrm{grace}}=1$ step is an administrative grace period. This implies that clusters crossing over at steps T1T-1 and $T$ have near-zero probability of observing $Y$, creating the design-induced near-boundary observation regime described in Section 1.
Cluster random effect.
The cluster-level random effect is drawn independently as $b_{j}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\sigma_{b}^{2})$ with $\sigma_{b}=0.184$ ($\sigma_{b}^{2}=0.034$, $\mathrm{ICC}\approx 0.05$), consistent with the ICC model of Section 2.3 and empirical ICC levels typical of public-health SW-CRTs (Hemming and Taljaard, 2020).
Baseline covariates.
Each individual has a $p=3$ dimensional baseline covariate vector:
$W_{ij}=(W_{ij,1},\,W_{ij,2},\,W_{ij,3})$
where $W_{ij,1}\sim\mathcal{N}(0,1)$, $W_{ij,2}\sim\mathrm{Bernoulli}(0.4)$, and $W_{ij,3}\sim\mathrm{Uniform}(0,1)$, all independent across individuals and clusters.
Secular time trend.
We specify the true secular time trend as a non-linear function of the calendar step to represent the realistic scenario where a linear time trend is misspecified:
$\lambda(t)=0.5\sin\!\left(\frac{\pi(t-1)}{T-1}\right)+0.3\left(\frac{t-1}{T-1}\right)^{2}.$
This ensures that parametric estimators using a simple linear-in-$t$ fixed effect will exhibit systematic secular time confounding bias.
Short-term surrogate.
The surrogate is generated from a linear model with a cluster random effect, treatment effect, covariate effects, and secular time trend:
$S_{ijt}$ $=0.8\,A_{jt}+0.4\,W_{ij,1}-0.3\,W_{ij,2}+0.5\,\lambda(t)+0.6\,b_{j}+\varepsilon^{S}_{ijt},$
$\varepsilon^{S}_{ijt}$ $\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\,0.5^{2}).$ (48)
The coefficient $0.6$ on $b_{j}$ induces a strong within-cluster surrogate–outcome correlation, making $S$ a useful bridge for the missing delayed outcome.
Observation indicator.
Administrative censoring follows a logistic model consistent with Assumption 3:
$\operatorname{logit}\bigl(\mathbb{P}(\Delta_{ijt}=1\mid S_{ijt},A_{jt},W_{ij},t)\bigr)$ $=3.0-2.5\cdot\mathbf{1}(t\geq T-1)-0.8\cdot\mathbf{1}(t=T)$
$\quad+0.4\,S_{ijt}+0.3\,A_{jt}-0.2\,W_{ij,3}.$ (49)
The large negative late-step coefficients concentrate censoring in the final two steps (overall rate $\approx 28\%$, rising to $\approx 80\%$ at $t=T$), creating the near-boundary observation regime for late-crossing clusters.
Delayed primary outcome.
Following the structural equation (6) and the ICC decomposition of Section 2.3:
$Y_{ijt}$ $=\Psi_{0}\,A_{jt}+0.5\,S_{ijt}+0.4\,W_{ij,1}-0.2\,W_{ij,2}+0.6\,W_{ij,3}+0.8\,\lambda(t)+b_{j}+\varepsilon^{Y}_{ijt},$
$\varepsilon^{Y}_{ijt}$ $\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\,0.8^{2}),$ (50)
where $\Psi_{0}=-0.28$ is the direct-effect coefficient, $\mathrm{ICC}=0.034/(0.034+0.64)\approx 0.05$.
True ATE.
The true total causal ATE is $\Psi^{*}=-0.28+0.5\times 0.8=0.12$ (direct effect plus mediated path through $S$), placed near the detection boundary for $J\approx 30$.
6.2 Simulation Scenarios
We organise the simulation study into four factorial scenario blocks. Each scenario is replicated 1,000 times.
Block I: Baseline performance across cluster counts.
The DGP is fixed at the specification above. We vary $J\in\{10,20,30,50,100\}$ at the baseline DGP ($\Psi^{*}=0.12$). This block assesses how finite-sample performance scales with the number of randomised clusters and whether the small-$J$ Lindeberg–Feller approximation is adequate for the cluster-level CLT.
Block II: Misspecified nuisance models.
$J=30$, $\Psi^{*}=0.12$. We consider three misspecification scenarios designed to probe the double robustness property of Theorem 2:
  1. (i)
    Outcome model misspecified, propensities correctly specified. The outcome models $\bar{Q}_{Y}$ and $\bar{Q}_{\mathrm{int}}$ are fit using a main-effects linear regression that omits the non-linear time trend $\lambda(t)$ and the interaction $S_{ijt}\times A_{jt}$; $g_{\Delta}$ is estimated by the correctly specified logistic model. Under the asymptotic double-robustness theory, consistency is still possible when the propensity side is correctly specified, but finite-sample performance may deteriorate if the misspecified outcome model leaves a large structured remainder.
  2. (ii)
    Propensities misspecified, outcome models correctly specified. $g_{\Delta}$ is estimated by logistic regression omitting $S_{ijt}$ (violating the surrogate-conditional independence structure of Assumption 3); $\bar{Q}_{Y}$ and $\bar{Q}_{\mathrm{int}}$ are correctly specified. Because $g_{A}$ is design-based and used as a plug-in (Remark 3), the treatment propensity remains correctly specified; only $g_{\Delta}$ is misspecified. Under the asymptotic double-robustness theory, the estimator is expected to remain stable when the outcome regressions are correctly specified, even if $g_{\Delta}$ is misspecified.
  3. (iii)
    Both nuisance models misspecified. Both the outcome models and $g_{\Delta}$ are misspecified as in (i) and (ii). This scenario is expected to produce estimator bias and is included as a negative control to verify that double robustness does not extend to simultaneous misspecification of both sides.
Block III: Increasing administrative censoring severity.
$J=30$, $\Psi^{*}=0.12$. We vary the administrative grace period $t_{\mathrm{grace}}\in\{0,1,2,3\}$, with $t_{\mathrm{grace}}=0$ representing strict administrative censoring (no grace period, so all clusters crossing over at the final two steps have $\Delta_{ijt}=0$ with probability approaching 1) and $t_{\mathrm{grace}}=3$ representing mild censoring. To ensure four distinct structural censoring regimes, we set $t_{\mathrm{lag}}=3$ for this block (versus the default $t_{\mathrm{lag}}=2$ in Blocks I, II, and IV). The overall censoring rates under the four values are approximately $\{43\%,28\%,16\%,8\%\}$. This block directly tests the theoretical prediction that IPCW variance explodes at high censoring rates while $\hat{\Psi}_{\mathrm{TMLE}}$ remains stable.
We compare three estimators: (i) GLMM (lme4, Gaussian family, linear-in-$t$ secular trend, cluster random intercept — deliberately misspecified in Blocks I and III–IV); (ii) IPCW, which upweights observed outcomes by the inverse estimated censoring propensity, with cluster-robust sandwich variance; and (iii) CV-TMLE (Supplement, Section S3, $V=10$ cluster-level folds), the recommended estimator. The treatment propensity $g_{A}$ is plugged in from the design for the TMLE variant.
6.3 Results
Performance is assessed over 1,000 replicates per scenario via bias, RMSE, 95% CI coverage, and power. Monte Carlo standard errors for coverage are $\approx 0.007$, so the nominal tolerance band is $[0.936,0.964]$.
Block I: Baseline performance across cluster counts.
Table 3 reports performance metrics for all three estimators across $J\in\{10,20,30,50,100\}$. The CV-TMLE achieves near-zero bias (<0.004<0.004) at all cluster counts. The GLMM exhibits persistent bias of approximately +0.07+0.07 at $J=30$ due to the misspecified secular time trend, and its coverage collapses from 0.52 at $J=10$ to 0.07 at $J=100$ as the standard error shrinks around the biased point estimate. IPCW exhibits substantial positive bias (approximately +0.22+0.22 at $J=30$) and markedly higher variance than the CV-TMLE, consistent with the theoretical prediction that inverse weighting becomes unstable when $g_{\Delta}$ concentrates near the positivity boundary for late-crossing clusters.
Coverage for CV-TMLE stabilises near 0.87–0.91 across $J\geq 20$, with no systematic improvement as $J$ grows to 100. Table 5 decomposes this gap: the sandwich estimates $\mathrm{Var}(\mathrm{EIC}_{j})/J$ but does not capture the variance of the second-order remainder $R_{2}$. The empirical variance exceeds the mean sandwich by a factor of about 1.6 at all $J$; this magnitude is broadly consistent with the observed coverage shortfall in the 0.87–0.89 range (see predicted column). This ratio is approximately constant because both terms are $O(J^{-1})$ under parametric nuisance rates. The theory suggests that this gap should diminish once nuisance estimation is sufficiently faster than the product-rate boundary; the present simulations indicate that the parametric nuisance specifications used here remain in a regime where that improvement is not yet visible. CV-TMLE power scales from 0.41 at $J=10$ to 1.00 at $J=100$, reflecting strong signal recovery together with very small empirical bias.
Table 2: Variance decomposition for CV-TMLE coverage (Block I). $\widehat{\mathrm{Var}}_{\mathrm{emp}}$: empirical variance of $\hat{\Psi}$ across 1,000 replicates. $\overline{\mathrm{Sand.}}$: mean sandwich variance estimate. Ratio >1>1 indicates sandwich underestimation attributable to unaccounted $R_{2}$ variance. Predicted coverage computed as $P(|Z|<t_{J-1}/\sqrt{\mathrm{Ratio}})$; agreement with actual coverage is approximate.
$J$ $\widehat{\mathrm{Var}}_{\mathrm{emp}}$ ($\times 10^{-3}$) $\overline{\mathrm{Sand.}}$ ($\times 10^{-3}$) Ratio Predicted Cov. Actual Cov.
10 8.44 4.80 1.76 0.912 0.878
20 4.83 2.83 1.71 0.891 0.869
30 2.93 1.84 1.59 0.895 0.892
50 1.80 1.11 1.62 0.886 0.866
100 0.87 0.54 1.61 0.883 0.878
Table 3: Block I simulation results: performance across cluster counts ($\Psi^{*}=0.12$, $T=7$, $n_{j}=40$, 1,000 replicates). “Cov.” denotes 95% CI coverage ($t_{J-1}$ critical values).
$J$ Estimator Bias Var $(\times 10^{-2})$ RMSE Cov. Power
10 GLMM +0.121+0.121 0.41 0.137 0.519 0.970
IPCW +0.235+0.235 0.61 0.248 0.985 0.172
CV-TMLE +0.004+0.004 0.84 0.092 0.878 0.410
20 GLMM +0.104+0.104 0.21 0.114 0.373 0.998
IPCW +0.231+0.231 0.33 0.238 0.934 0.649
CV-TMLE +0.001+0.001 0.48 0.070 0.869 0.554
30 GLMM +0.066+0.066 0.16 0.077 0.597 0.999
IPCW +0.219+0.219 0.22 0.224 0.787 0.939
CV-TMLE +0.000+0.000 0.29 0.054 0.892 0.727
50 GLMM +0.083+0.083 0.08 0.088 0.203 1.000
IPCW +0.224+0.224 0.12 0.227 0.220 1.000
CV-TMLE +0.003+0.003 0.18 0.043 0.866 0.897
100 GLMM +0.074+0.074 0.05 0.077 0.068 1.000
IPCW +0.220+0.220 0.06 0.221 0.000 1.000
CV-TMLE +0.000+0.000 0.09 0.030 0.878 0.997
Block II: Misspecified nuisance models.
Table 4 reports performance at $J=30$ under the three misspecification regimes of Section 6.2. Under scenario (ii) (outcome models correct, $g_{\Delta}$ misspecified), CV-TMLE achieves near-zero bias (<0.001<0.001) and coverage of 0.921, providing supportive finite-sample evidence for the asymptotic double-robustness result in the regime where the outcome regressions are correctly specified: the correctly specified outcome models drive the product-rate condition (42) toward $o_{P}(J^{-1/2})$, consistent with Remark 5.
Under scenario (i) (outcome models misspecified, $g_{\Delta}$ correct), CV-TMLE exhibits bias of 0.100 with coverage 0.426. This illustrates an important finite-sample limitation of the asymptotic double-robustness guarantee: when the misspecified outcome regression leaves a large structured error, correct specification of $g_{\Delta}$ alone may not be sufficient to deliver good finite-sample performance at $J=30$. In this setting, the targeting step appears to reduce but not eliminate the residual bias induced by the misspecified outcome model.
Under scenario (iii) (both sides misspecified), CV-TMLE is biased ($\mathrm{Bias}=0.092$) and under-covers (coverage 0.593), as expected when both nuisance components are misspecified, illustrating that the asymptotic double-robustness guarantee does not extend to simultaneous misspecification. The GLMM is biased identically in all three scenarios (bias =0.066=0.066, coverage =0.597=0.597) because it uses neither $g_{\Delta}$ nor the surrogate bridge. IPCW is biased in all scenarios due to the near-boundary observation probabilities for late-crossing clusters, with identical results in scenarios (ii) and (iii) where the same $g_{\Delta}$ misspecification applies.
Table 4: Block II simulation results: double robustness under nuisance misspecification ($J=30$, $\Psi^{*}=0.12$, 1,000 replicates). Scenario labels: (i) outcome models only misspecified; (ii) censoring propensity only misspecified; (iii) both misspecified.
Scenario Estimator Bias RMSE Coverage Power
(i): $\bar{Q}_{Y},\bar{Q}_{\mathrm{int}}$ misspec. GLMM +0.066+0.066 0.077 0.597 0.999
IPCW +0.219+0.219 0.224 0.787 0.939
CV-TMLE +0.100+0.100 0.113 0.426 0.990
(ii): $g_{\Delta}$ misspec. GLMM +0.066+0.066 0.077 0.597 0.999
IPCW +0.131+0.131 0.142 0.996 0.156
CV-TMLE +0.000+0.000 0.057 0.921 0.633
(iii): Both misspec. GLMM +0.066+0.066 0.077 0.597 0.999
IPCW +0.131+0.131 0.142 0.996 0.156
CV-TMLE +0.092+0.092 0.106 0.593 0.965
Block III: Bias amplification under increasing censoring.
Figure 2 displays estimator performance as censoring increases from 8% to 43%. Across the censoring range, CV-TMLE maintains substantially smaller bias than IPCW and GLMM, while its coverage declines from 0.92 to 0.77 as censoring becomes more severe (Panels A–B). This pattern is qualitatively consistent with the remainder-variance explanation from Block I: the point estimator remains comparatively stable, but the sandwich interval becomes increasingly incomplete as censoring amplifies the contribution of nuisance-estimation error. GLMM and IPCW coverage collapses to near zero. The near-100% rejection rates of GLMM and IPCW at high censoring (Panel C) are driven largely by bias rather than accurate signal recovery, as reflected in their simultaneous near-zero coverage. CV-TMLE power (0.65–0.73) reflects genuine signal detection.
Figure 2: Block III: bias (A), coverage (B), and power (C) under increasing censoring ($J=30$, $t_{\mathrm{lag}}=3$, 1,000 replicates). CV-TMLE bias stays near zero while GLMM/IPCW bias grows to 0.32. High GLMM/IPCW rejection rates at heavy censoring reflect bias, not signal (near-zero coverage).
ICC heterogeneity (Block IV).
Additional simulations varying $\sigma_{b}^{2}\in\{0.006,0.034,0.071,0.160\}$ (ICC $\approx$ 0.01–0.20) confirm that the cluster-robust sandwich estimator maintains valid coverage across the empirically observed ICC range in implementation-science SW-CRTs; full results are in Supplement, Section S4.
Table 5: Variance decomposition for CV-TMLE coverage (Block I). $\widehat{\mathrm{Var}}_{\mathrm{emp}}$: empirical variance of $\hat{\Psi}$ across 1,000 replicates. $\overline{\mathrm{Sand.}}$: mean sandwich variance estimate. Ratio >1>1 indicates sandwich underestimation attributable to unaccounted $R_{2}$ variance. Predicted coverage computed as $P(|Z|<t_{J-1}/\sqrt{\mathrm{Ratio}})$; agreement with actual coverage is approximate.
$J$ $\widehat{\mathrm{Var}}_{\mathrm{emp}}$ ($\times 10^{-3}$) $\overline{\mathrm{Sand.}}$ ($\times 10^{-3}$) Ratio Predicted Cov. Actual Cov.
10 8.44 4.80 1.76 0.912 0.878
20 4.83 2.83 1.71 0.891 0.869
30 2.93 1.84 1.59 0.895 0.892
50 1.80 1.11 1.62 0.886 0.866
100 0.87 0.54 1.61 0.883 0.878
7 Design-Calibrated Illustration: Washington State EPT Trial
Setting.
The EPT trial (Golden et al., 2015) assessed whether a public health programme promoting free patient-delivered partner therapy could reduce chlamydia positivity among women aged 14–25 in Washington State. The four-wave stepped-wedge design randomized $J=23$ local health jurisdictions (LHJs) across $T=5$ time periods (one baseline, four post-baseline steps of six to eight months each): six LHJs crossed over at each of waves 1–3 and five at wave 4, with approximately 162 chlamydia tests per LHJ per period at sentinel clinical sites. The delayed primary outcome is 12-month chlamydia test positivity. This is administratively censored for late-crossing LHJs: wave-4 LHJs ($t_{j}=4$) have under 12 months of post-crossover follow-up, yielding an administrative censoring rate of 86% in that wave; wave-3 LHJs are partially censored (40%) from period 3 onward; overall 33.7% of records have missing 12-month outcomes. EPT uptake — whether the index patient accepted patient-delivered partner therapy within 30 days of STI diagnosis — is complete for all subjects regardless of wave and satisfies the three surrogate criteria of Remark 2: causal proximity to the intervention, temporal availability before the 12-month outcome, and biological stationarity.
We generate a replication dataset from the SCM of Section 2, with parameters fixed to match the published summary statistics of Golden et al. (2015): control positivity 5.0%, ICC $\approx 0.05$ (cluster-level CV =0.0165=0.0165), N=18,630N=18{,}630 individual records, and a treatment effect of 10% relative reduction on the log-risk scale. The oracle ATE is therefore known by construction: $\Psi_{0}=-0.0039$ (risk difference), consistent with the original trial’s null finding. All calibration code is in the software repository.
Results.
All three estimators produce point estimates within 0.003 of the oracle ATE $\Psi_{0}=-0.0039$ and achieve coverage (Table 6; Figure 3). The substantive comparison is CI width: IPCW (width 0.068) is twice that of SA-TMLE (0.034), reflecting variance inflation from near-zero wave-4 censoring weights — precisely the near-boundary regime characterised by Proposition 2. The GLMM achieves the narrowest CI (0.026) at the cost of requiring correct model specification for validity.
Table 6: Oracle comparison for the design-calibrated EPT illustration. Data generated under the SCM of Section 2 with parameters matching Golden et al. (2015): $J=23$ clusters, $T=5$ periods, overall censoring 33.7%, wave-4 censoring 86.5%, oracle ATE =0.0039=-0.0039. The exercise assesses which estimator’s CI covers the known truth and at what cost in bias and CI width; it does not constitute an inferential claim about the EPT intervention.
Estimator ATE SE 95% CI Oracle covered?
Oracle ($\Psi_{0}$, known by construction) 0.0039-0.0039
GLMM (complete-case) 0.0012-0.0012 $0.0066$ (0.014, 0.012)(-0.014,\;0.012) Yes
IPCW 0.0059-0.0059 $0.0163$ (0.040, 0.028)(-0.040,\;0.028) Yes
SA-TMLE (proposed) 0.0008-0.0008 $0.0082$ (0.018, 0.016)(-0.018,\;0.016) Yes
Figure 3: Oracle comparison for the design-calibrated EPT illustration (calibrated to Golden et al. 2015). Horizontal bars are 95% cluster-robust confidence intervals; the vertical dashed line marks zero; the dotted line marks the known oracle $\Psi_{0}=-0.0039$. All three estimators cover the oracle, with point estimates within 0.003 of the truth. The key comparison is CI width: IPCW (width 0.068) is twice as wide as SA-TMLE (0.034), reflecting variance inflation from near-zero wave-4 censoring weights. GLMM achieves the narrowest CI (0.026) at the cost of model dependence. This figure illustrates oracle coverage under a calibrated design, not the causal effect of EPT.
8 Discussion
We proposed a Surrogate-Assisted TMLE linking the longitudinal G-computation formula (1) to the cluster-level EIC of Section 3, providing an asymptotically doubly robust, non-parametric solution to the administrative censoring of late-crossing clusters in SW-CRTs. Standard IPCW fails in this setting because the censoring propensity $g_{\Delta}$ can become arbitrarily small for late-crossing clusters (Remark 3), producing highly unstable weights. The Surrogate-Assisted TMLE bypasses this instability by bridging missing outcomes through $S$, which is observed for all individuals regardless of the censoring horizon; the simulation study of Section 6 confirms near-zero bias and markedly lower variance than IPCW across all scenarios evaluated.
Finite-sample coverage.
The observed CV-TMLE coverage of 0.87–0.91 highlights a finite-sample regime in which the asymptotic approximation is directionally informative but not yet fully calibrated. Our finite-sample experiments should be read as showing qualitative agreement with the asymptotic theory rather than exact small-sample calibration; in the regimes studied here, the remaining undercoverage appears driven by nuisance-estimation error that is not yet negligible at the available cluster counts. Table 5 documents that the empirical variance exceeds the mean sandwich by a factor of $\approx 1.6$ at all $J$, qualitatively consistent with the boundary case $\alpha+\beta\approx 1/2$ of Corollary S2 (Supplement, Section S2.7). The $t_{J-1}$ interval is a practical safeguard, but not a full cure; practical remedies include a richer Super Learner library, a cluster bootstrap, or the variance inflation correction described in Supplement, Section S2.7.
An important methodological implication of Proposition 2 is that the nested G-computation formula (1) cannot be targeted by a straightforward DML one-step without an additional rate condition on the conditional surrogate density $f_{S}$. The SA-TMLE’s two-stage nested fluctuation eliminates this requirement by constructing $\hat{\bar{Q}}^{*}_{\mathrm{int}}$ as a propensity-weighted empirical mean of $\hat{\bar{Q}}^{*}_{Y}$, marginalizing over the observed surrogate distribution without estimating $f_{S}$.
Assumption diagnostics and sensitivity.
The validity of the estimator hinges on Assumption 3 (Surrogate-Mediated MAR), whose plausibility is discussed in Remark 2. Assumption 3 is partially testable on the fully-observed sub-cohort via the augmented regression:
$\mathbb{E}[Y\mid S,A,W,t,\Delta=1]\;=\;\mathbb{E}[Y\mid S,A,W,t,\hat{r}(A,W,t),\Delta=1],$ (51)
where $\hat{r}(A,W,t)=\hat{P}(\Delta=1\mid A,W,t)/\hat{P}(\Delta=1\mid S,A,W,t)$; a zero coefficient on $\hat{r}$ is consistent with Assumption 3.
Sensitivity analysis under MAR violation.
When Assumption 3 fails, the bias of $\hat{\Psi}_{\mathrm{TMLE}}$ can be characterized via a tipping-point analysis. Let $\gamma(S,A,W,t)$ denote the magnitude of the MAR violation:
$\gamma(S,A,W,t)\;\equiv\;\mathbb{E}[Y\mid S,A,W,t,\Delta=0]\;-\;\mathbb{E}[Y\mid S,A,W,t,\Delta=1].$ (52)
This is the conditional mean difference in $Y$ between the unobserved ($\Delta=0$) and observed ($\Delta=1$) sub-cohorts, given the surrogate. Under Assumption 3, $\gamma\equiv 0$. Under a violation of magnitude $\gamma$, the leading-order bias of $\hat{\Psi}_{\mathrm{TMLE}}$ is:
$\mathrm{Bias}(\hat{\Psi}_{\mathrm{TMLE}})\;\approx\;\mathbb{E}\!\left[\frac{I(A=1)(1-g_{\Delta})}{g_{\Delta}}\,\gamma(S,1,W,t)\;-\;\frac{I(A=0)(1-g_{\Delta})}{g_{\Delta}}\,\gamma(S,0,W,t)\right],$ (53)
where the expectation is over the marginal distribution and $1-g_{\Delta}$ is the censoring probability. Equation (53) reveals two sensitivity drivers: the censoring rate $(1-g_{\Delta})/g_{\Delta}$ (large for late-crossing clusters) and the differential violation $\gamma(S,1,W,t)-\gamma(S,0,W,t)$ across arms. The tipping-point $\gamma^{*}$ is obtained by solving $\hat{\Psi}_{\mathrm{TMLE}}=\hat{\mathrm{Bias}}(\gamma^{*})$; values implausibly large given subject-matter knowledge indicate a robust conclusion. This extends Scharfstein et al. (1999) to the surrogate-mediated setting and is analogous to the $E$-value of VanderWeele and Ding (2017); we recommend reporting $\gamma^{*}$ routinely.
More broadly, the results here suggest that surrogate-assisted targeted learning can serve as a template for a wider class of nested semiparametric functionals in which a primary outcome is observed sparsely but an intermediate variable remains broadly available. Although the bridge formulation extends beyond stepped-wedge trials, the present theoretical and numerical development is calibrated to the stepped-wedge setting; extension to other longitudinal designs is a promising but separate task. The key lesson is not only that surrogate information can stabilize inference under administrative censoring, but that the geometry of the nested bridge functional itself may require estimator constructions beyond standard one-step debiasing. Extending these ideas to multistage treatments, informative surrogate measurement, and more general longitudinal bridge structures is a natural direction for future work.
8.1 Future Directions
Two directions remain: (i) extending the cluster-level EIC to time-to-event outcomes with continuous administrative censoring; and (ii) establishing non-asymptotic coverage bounds for the CV-TMLE when nuisance classes are unbounded, which requires bracketing entropy bounds beyond those currently available for Super Learner ensembles. The $t_{J-1}$ confidence interval (41) provides a practical safeguard, formally justified by Theorem S1.
Supplementary Material
The Supplementary Material contains: Section S1, the Berry–Esseen coverage bound (Theorem S1 and Corollary S1); Section S2, the derivation of the efficient influence curve including the tangent space factorization, total EIC, cluster-robust variance estimator, second-order remainder and double robustness proof, and the DML cross-product remainder (Proposition 1 proof); Section S3, the Super Learner specification and cluster-level CV-TMLE implementation details.
All analyses were conducted in R (version 4.3 or later). The Surrogate-Assisted CV-TMLE estimator, including the cluster-level Super Learner wrapper, the nested fluctuation step, the cluster-robust sandwich variance estimator, and all four simulation scenario generators (Blocks I–III of Section 6, and the Block IV ICC simulations), are implemented in the open-source R package swcrtSurrTMLE, available at https://github.com/[author]/swcrtSurrTMLE. The package depends on SuperLearner (Polley and van der Laan, 2010) for ensemble estimation, lme4 for the GLMM comparator, and tmle for reference TMLE implementations. A fully reproducible script that replicates all tables and figures in this paper is included in the package vignette and archived on Zenodo at https://doi.org/10.5281/zenodo.[XXXXX] (DOI to be assigned upon acceptance).

### 15 Mar 2026 v2

@@
-Surrogate-Assisted Targeted Learning for Delayed Outcomes under Administrative Censoring
+Surrogate-Assisted Targeted Learning for Nested Bridge Functionals under Administrative Censoring
 Abstract
-Delayed primary outcomes and administratively censored follow-up create a general semiparametric estimation problem: the target causal functional depends on an endpoint observed only for a shrinking subset of units at analysis time, while earlier surrogate measurements remain widely available.
-In such settings, inverse-probability-weighted estimators can become unstable as observation probabilities approach the positivity boundary, and complete-case model-based analyses can be highly sensitive to outcome-model specification.
-We develop a surrogate-assisted targeted minimum loss estimator for this nested causal functional.
-Identification proceeds through a surrogate-bridge representation that integrates an observed-outcome regression over the conditional surrogate distribution, thereby avoiding inverse observation weights in the target parameter itself.
-We show that the estimator is asymptotically linear and doubly robust (in the sense that first-order bias vanishes when either nuisance component is consistently estimated), and we characterize two structural features of the problem: under surrogate-mediated missing at random, the censoring mechanism contributes no separate tangent-space component to the efficient influence function; and for nested bridge functionals, a one-step debiased machine-learning construction leaves a second-order cross-product remainder involving the conditional surrogate law.
-The proposed two-stage targeting step removes this term without requiring direct estimation of that law.
-Simulation studies demonstrate stable finite-sample performance under substantial administrative censoring, and a design-calibrated analysis based on the Washington State EPT study illustrates the method in a realistic stepped-wedge cluster-randomized setting.
+We study semiparametric estimation of the average treatment effect when the target functional belongs to the nested bridge class: identification requires integrating an outcome regression over a conditional intermediate distribution rather than inverse-weighting incomplete observations.
+This structure arises whenever a primary outcome is administratively censored for a non-negligible fraction of units while an earlier surrogate remains broadly available, and it generates a distinctive obstruction for standard first-order estimators: a one-step debiased machine-learning construction leaves a second-order cross-product remainder $R_{SY}$ involving the conditional surrogate law $f_{S}$ that cross-fitting does not eliminate and that has no doubly-robust complement in the efficient influence function.
+We develop a surrogate-assisted targeted minimum loss estimator (SA-TMLE) that resolves this obstruction via a two-stage nested fluctuation step, and establish three structural results: (i) under surrogate-mediated missing at random, the censoring mechanism contributes no separate tangent-space component to the efficient influence function; (ii) the standard DML one-step remainder contains $R_{SY}$ as an irreducible second-order term (Proposition 1); and (iii) the two-stage targeting step absorbs $R_{SY}$ into the efficient score without estimating $f_{S}$, yielding an asymptotically linear, doubly robust estimator under a single product-rate condition (Theorem 2).
+For clustered data, valid inference requires summation rather than averaging of individual influence contributions at the cluster level (Lemma 2); a leave-one-cluster-out jackknife applied to the full estimator restores near-nominal finite-sample coverage where the first-order sandwich interval undercovers.
+Stepped-wedge cluster-randomized trials, in which calendar-driven rollout creates near-boundary observation probabilities for late-crossing clusters, provide the sharpest instance of this functional class and serve as the running example.
+Simulations confirm near-zero bias and jackknife coverage of 0.956–0.968 across all cluster sizes evaluated, against a backdrop where inverse-weighted competitors break down entirely at high censoring rates.
 1 Introduction
-Many modern studies measure the primary endpoint only after substantial delay, while shorter-term surrogate outcomes are available much earlier in follow-up.
-When analysis occurs before all primary outcomes mature, the resulting observed-data structure poses a distinctive semiparametric problem: the estimand depends on a delayed endpoint observed only for a subset of units whose observation probability may approach the positivity boundary, yet a surrogate remains broadly observed across the design space.
-In such settings, estimators that rely directly on inverse observation weights can become unstable as observation probabilities become small, whereas complete-case model-based analyses may inherit strong sensitivity to outcome-model specification.
-This paper studies that problem through a surrogate-bridge formulation for causal effect estimation under administrative censoring.
-The key idea is to identify the target functional — the average treatment effect $\Psi(P_{0})=\mathbb{E}[Y(1)-Y(0)]$ — by integrating an observed-outcome regression over the conditional surrogate law, rather than by placing the inverse observation probability directly in the target parameter.
-This perspective leads naturally to a targeted learning estimator and exposes new structure in the corresponding semiparametric theory.
-A stepped-wedge cluster randomized trial (SW-CRT) provides a particularly sharp motivating example.
-Because treatment rollout is tied to calendar time, late-crossing clusters may have little opportunity to realize the delayed endpoint before study closure.
-But the underlying statistical issue is broader than stepped-wedge designs: whenever an early surrogate is observed widely and a primary endpoint is administratively censored at analysis time, the analyst faces the same nested surrogate-bridge problem.
-We therefore use the stepped-wedge setting as a concrete and practically important instance of a more general semiparametric estimation problem.
+A class of causal functionals in modern longitudinal studies shares the following semiparametric structure.
+The target parameter — an average treatment effect — is identified not by inverse-weighting incomplete observations, but by integrating an outcome regression over the conditional distribution of an intermediate variable.
+We call this a nested bridge functional: its efficient score has a cross-product structure that creates a specific failure mode for standard first-order estimators, and resolving that failure mode requires a targeting step that goes beyond ordinary debiased machine learning.
+The observational setting that generates this structure is one where a primary endpoint requires substantial follow-up to realize, so that only units enrolled early enough have a fully matured observation by the time of analysis.
+A shorter-term surrogate is observed broadly across the study population regardless of enrollment timing.
+Rather than inverse-weighting the incomplete observations by $g_{\Delta}^{-1}$ — which becomes unstable whenever $g_{\Delta}$ concentrates near zero — one identifies the average treatment effect by integrating the complete-case outcome regression $\bar{Q}_{Y}(S,a,W,t)\equiv\mathbb{E}[Y\mid S,A=a,W,t,\Delta=1]$ over the treatment-specific conditional distribution of the surrogate:
+$\Psi(P_{0})\;=\;\mathbb{E}_{W,t}\Bigl(\mathbb{E}_{S\mid A=1,W,t}\bigl[\bar{Q}_{Y}(S,1,W,t)\bigr]-\mathbb{E}_{S\mid A=0,W,t}\bigl[\bar{Q}_{Y}(S,0,W,t)\bigr]\Bigr).$ (1)
+This representation requires only that $\bar{Q}_{Y}$ be estimable on the observed sub-cohort $\{\Delta=1\}$, not that $g_{\Delta}$ be bounded away from zero.
+The price is a new semiparametric complication: the functional (1) is nested in the sense that the outer integration is over the conditional surrogate law $f_{S}$, and the estimator must marginalize over $f_{S}$ without placing it in a denominator.
+The core semiparametric obstruction.
+The nested structure of (1) creates a specific failure mode for standard debiased machine-learning (DML) estimators.
+A natural one-step construction estimates $\bar{Q}_{\mathrm{int}}(a,W,t)\equiv\int\bar{Q}_{Y}(s,a,W,t)\,f_{S}(s\mid a,W,t)\,d\mu(s)$ by plugging in $\hat{\bar{Q}}_{Y}$ and $\hat{f}_{S}$ separately, leaving a second-order cross-product remainder
+$R_{SY}\;=\;\sum_{a}\iint\bigl(\hat{\bar{Q}}_{Y}-\bar{Q}^{0}_{Y}\bigr)(s,a,w,t)\,\bigl(\hat{f}_{S}-f^{0}_{S}\bigr)(s\mid a,w,t)\;d\mu(s)\,dP_{0}(w,t).$ (2)
+Unlike the standard doubly-robust remainder, $R_{SY}$ has no complement in the efficient influence function: no nuisance parameter whose consistent estimation forces $R_{SY}=0$.
+DML cross-fitting eliminates first-order empirical-process terms but leaves $R_{SY}$ unchanged, since $R_{SY}$ is a $P_{0}$-expectation of a second-order product.
+Controlling $R_{SY}=o_{P}(J^{-1/2})$ therefore requires either a rate condition on $\hat{f}_{S}$ or an estimator that avoids placing $\hat{f}_{S}$ in the expansion.
+The SA-TMLE takes the second path: its two-stage nested fluctuation enforces $N^{-1}\sum D_{S,a}(O_{ijt})=0$, constraining $\hat{\bar{Q}}^{*}_{\mathrm{int}}$ to a propensity-weighted empirical mean of $\hat{\bar{Q}}^{*}_{Y}$ over observed surrogate values, absorbing $R_{SY}$ without ever estimating $f_{S}$.
+Motivating instance.
+Stepped-wedge cluster-randomized trials (SW-CRTs) provide a sharp instance of this structure.
+Calendar-determined treatment rollout means late-crossing clusters accumulate few primary-outcome observations before the administrative end date, creating a design-induced near-boundary regime $g_{\Delta}\to 0$ for those clusters.
+A short-term surrogate observed at crossover is available for all clusters, making the surrogate-bridge framework directly applicable.
+We use the SW-CRT as the running example throughout; the framework is not tied to this design.
 Limitations of standard approaches.
-Parametric mixed models (Hussey and Hughes, 2007; Hemming and Taljaard, 2020; Hughes et al., 2020) restrict attention to the complete-case sub-cohort $\{\Delta_{ijt}=1\}$ and require correct specification of both the secular time trend and the missing-data mechanism.
-IPCW estimators (Robins et al., 1994; Bang and Robins, 2005) place $g_{\Delta}^{-1}$ directly in the estimating equation; even moderate concentration of $g_{\Delta}$ near zero for late-crossing clusters produces severe variance inflation, and in the near-boundary regime these estimators may fail to admit regular asymptotic approximations.
+IPCW estimators (Robins et al., 1994; Bang and Robins, 2005) place $g_{\Delta}^{-1}$ directly in the estimating equation, producing severe variance inflation as $g_{\Delta}\to 0$ and potentially failing to admit regular asymptotic approximations near the boundary.
+Parametric mixed models (Hussey and Hughes, 2007; Hemming and Taljaard, 2020; Hughes et al., 2020) restrict attention to the complete-case sub-cohort and require correct specification of the secular time trend and censoring mechanism (Kasza et al., 2019).
 Contributions.
-This paper makes three methodological contributions.
-First, we identify a class of causal functionals under surrogate-mediated administrative censoring through a nested bridge representation (Theorem 1) that avoids inverse observation weights in the target parameter.
-The surrogate-bridge G-computation formula identifies $\Psi(P_{0})$ via nested conditional expectations over the observed surrogate distribution, replacing the unstable $g_{\Delta}^{-1}$ with a support positivity condition on the complete-case outcome regression.
-Second, we characterize the efficient-score structure of this problem (Section 3): under surrogate-mediated missing at random, the censoring mechanism contributes no separate tangent-space component to the efficient influence function (Lemma 1), so estimating $g_{\Delta}$ does not reduce the efficiency bound; and under clustered dependence, valid inference requires aggregation of influence contributions at the cluster level rather than averaging (Lemma 2).
-Third, we show that for this nested functional, a standard one-step debiased machine-learning construction generally leaves a second-order cross-product remainder $R_{SY}$ involving the conditional surrogate law $f_{S}$ (Proposition 2); the proposed two-stage targeted minimum loss procedure removes that term without requiring direct estimation of $f_{S}$, achieving $\sqrt{J}$-consistency under a product-rate condition on the remaining nuisance estimators (Theorem 2).
+This paper makes three contributions to semiparametric estimation theory, instantiated in the SW-CRT setting.
+First, we identify the nested bridge functional class under surrogate-mediated administrative censoring and establish a G-computation representation (Theorem 1) that avoids inverse observation weights entirely.
+The representation replaces the unstable $g_{\Delta}^{-1}$ with a support positivity condition on the complete-case outcome regression, extending the scope of doubly robust estimation to near-boundary censoring regimes.
+Second, we characterize the efficient-score geometry of this functional class (Section 3).
+Three structural results follow: the censoring mechanism contributes no separate tangent-space component under surrogate-mediated MAR (Lemma 1); the nested cross-product remainder $R_{SY}$ is an irreducible second-order obstruction with no doubly-robust complement (Proposition 1); and valid inference under clustered dependence requires summation rather than averaging of individual influence contributions at the cluster level (Lemma 2).
+Third, we construct the SA-TMLE and prove that its two-stage nested fluctuation absorbs $R_{SY}$ without estimating $f_{S}$, achieving $\sqrt{J}$-consistent asymptotic linearity under a single product-rate condition (Theorem 2).
+For finite-sample inference, a leave-one-cluster-out jackknife applied to the full estimator captures the fluctuation-step variance that the first-order sandwich misses, restoring near-nominal coverage across the cluster-count range evaluated.
 Related statistical work.
-Our work sits at the intersection of semiparametric missing-data theory, targeted learning, and modern debiased machine learning.
-Classical AIPW theory (Robins et al., 1994; Bickel et al., 1993) characterizes efficient estimation with missing outcomes via tangent space decomposition, but does not address the nested bridge structure considered here.
-Existing SW-CRT methodology (Hussey and Hughes, 2007; Hemming and Taljaard, 2020; Kasza et al., 2019; Hughes et al., 2020) largely remains centered on parametric mixed models and does not accommodate the present surrogate-assisted functional.
-The surrogate outcome literature (Gilbert et al., 2008; Frangakis and Rubin, 2002) addresses a complementary problem — evaluating treatment effects when the primary outcome is rare, not delayed — and their identification strategies do not extend to the near-boundary censoring regime.
+Classical AIPW theory (Robins et al., 1994; Bickel et al., 1993) characterizes efficient estimation with missing outcomes but does not address the nested bridge structure or the $R_{SY}$ obstruction.
+The surrogate outcome literature (Gilbert et al., 2008; Frangakis and Rubin, 2002) addresses the complementary problem of rare (not delayed) primary outcomes; their identification strategies do not extend to the near-boundary censoring regime.
 Luedtke et al.
-(2017) construct the closest prior estimator: a semiparametric surrogate-index estimator for i.i.d.
-data whose identification formula is structurally related to ours.
-However, their asymptotic linearity argument requires a uniform positivity lower bound on $g_{\Delta}$, which the SW-CRT’s calendar-driven censoring structure violates for late-crossing clusters; the present surrogate-bridge approach does not place $g_{\Delta}^{-1}$ in the target parameter and therefore does not require this condition.
-Cluster-level targeted learning (Balzer et al., 2016, 2021) establishes the importance of clusterwise influence aggregation for valid sandwich inference in group-randomized trials, but without the bridge-based identification and nested targeting problem studied here.
-The debiased machine learning framework (Chernozhukov et al., 2018) provides general product-rate guarantees via sample splitting, yet in the present setting a one-step construction leaves a nested cross-product remainder $R_{SY}$ involving the conditional surrogate law $f_{S}$ that cross-fitting does not eliminate.
-The contribution of this paper is to isolate that remainder explicitly (Proposition 2) and show how a second targeting step absorbs it without estimating $f_{S}$.
-[^1] See Kennedy (2022) for a comprehensive review of debiased machine learning.
-Proposition 2 gives the formal derivation.
+(2017) construct a semiparametric surrogate-index estimator for i.i.d.
+data structurally related to ours, but their asymptotic linearity requires a uniform positivity lower bound on $g_{\Delta}$ that the present approach does not need.
+The DML framework (Chernozhukov et al., 2018) provides product-rate guarantees via sample splitting, but cross-fitting does not eliminate $R_{SY}$; this paper isolates that remainder (Proposition 1) and resolves it via a second targeting step.
+[^1] See Kennedy (2022) for a review of debiased ML.
+Cluster-level TMLE (Balzer et al., 2016, 2021) establishes clusterwise influence aggregation for group-randomized trials without the nested bridge structure studied here.
 Section 2 defines the observed data structure and establishes non-parametric identification via the surrogate-bridge representation.
-Section 3 develops the semiparametric theory: the EIC decomposition, the vanishing censoring-mechanism component (Lemma 1), the cluster-level summation rule (Lemma 2), and the nested cross-product remainder that makes two-stage targeting necessary (Proposition 2).
-Section 4 constructs the SA-TMLE as the direct solution to the theoretical constraints identified in Section 3.
-Section 5 establishes asymptotic linearity, double robustness, and the Berry–Esseen finite-sample coverage bound.
-Section 6 reports Monte Carlo experiments across three scenario blocks.
-Section 7 provides a brief design-calibrated illustration using the Washington State EPT trial.
-Although our motivating application is a stepped-wedge cluster trial, the framework is not tied to that design.
-The same statistical structure arises whenever a delayed primary outcome is observed only for a subset of units at analysis time, an earlier surrogate is broadly available, and identification proceeds through a nested bridge representation.
-The paper is best understood not as a stepped-wedge method but as a general semiparametric framework for nested surrogate-bridge inference under administrative censoring, with stepped-wedge trials providing a particularly consequential use case.
-Table 1: When to use each estimator: a practical comparison for stepped-wedge trials with delayed outcomes.
-GLMM IPCW/AIPW SA-TMLE (proposed)
-Target parameter Conditional treatment effect (model-based) Population ATE Population ATE
-Key assumptions Correct outcome model specification (secular trend, random effects) Observation positivity: $g_{\Delta}^{-1}$ must be well-behaved Support positivity for $E[Y\mid S,A,W,t]$; surrogate-mediated MAR
-Handles heavy censoring? No: complete-case analysis drops late-crossing clusters entirely Poorly: variance inflates as $g_{\Delta}\to 0$ Yes: bypasses $g_{\Delta}^{-1}$ via surrogate bridge
-Requires surrogate? No No Yes
-Double robustness? No Yes (but requires bounded $g_{\Delta}^{-1}$) Yes (asymptotically; finite-sample behavior may be asymmetric under nuisance misspecification)
-Nonparametric nuisance? No (parametric outcome model) Yes (machine learning compatible) Yes (machine learning compatible)
-Best suited for Low censoring, well-understood secular trends, no surrogate available Moderate censoring with good observation-positivity overlap Substantial administrative censoring with an available early surrogate
+Section 3 develops the semiparametric theory, including the EIC decomposition, the vanishing censoring-mechanism component (Lemma 1), the cluster-level summation rule (Lemma 2), and Proposition 1 characterizing $R_{SY}$.
+Section 4 constructs the SA-TMLE.
+Section 5 establishes asymptotic linearity, double robustness, and the Berry–Esseen coverage bound.
+Section 6 reports Monte Carlo experiments.
+Section 7 provides the design-calibrated EPT illustration.
 2 Data Structure and Identification
-We first describe the observed-data structure and identification strategy in the stepped-wedge cluster-randomized setting, which serves as our running example of a broader class of surrogate-assisted problems with administratively censored delayed outcomes.
+We now formalize the observed-data structure and identification strategy for the general nested surrogate-bridge problem described in Section 1.
+The framework applies whenever a delayed primary outcome is administratively censored for a subset of units at analysis time and an earlier surrogate is broadly observed.
+We instantiate it throughout in the stepped-wedge cluster-randomized setting, which makes the positivity boundary concrete and provides the simulation and illustration target of later sections; the identification theorem and semiparametric theory of Sections 2.23 are stated for the general structure.
 Consider a SW-CRT with $J$ clusters, $T$ calendar time steps, and $n_{j}$ individuals in cluster $j$ ($N=\sum_{j}n_{j}$ total).
@@
 Three structural features deserve emphasis.
-First, $A_{jt}$ is deterministic given $\tau_{j}$; the treatment mechanism is known by design and does not require data-adaptive estimation (see Remark 4).
+First, $A_{jt}$ is deterministic given $\tau_{j}$; the treatment mechanism is known by design and $g_{A}(1\mid W,t)=(t-1)/(T-1)$ is plugged in directly rather than estimated.
 Second, the cluster random effect $b_{j}$ enters both $f_{S}$ and $f_{Y}$, inducing the ICC among all observations within cluster $j$ (Remark 1).
@@
 Remark 1 (Intra-cluster correlation).
-The shared cluster random effect $b_{j}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\sigma^{2}_{b})$ enters both $f_{S}$ and $f_{Y}$, inducing ICC $\sigma^{2}_{b}/(\sigma^{2}_{b}+\sigma^{2}_{\varepsilon})$ among individuals within a cluster.
-The two methodological consequences are: (i) inference must use the cluster-robust sandwich estimator of Section 3; and (ii) Super Learner cross-validation must be performed at the cluster level (Supplement, Section S3).
-The TMLE estimator itself makes no parametric assumption about $b_{j}$; the Gaussian model governs the simulation DGP only.
+The shared $b_{j}$ induces ICC $\sigma^{2}_{b}/(\sigma^{2}_{b}+\sigma^{2}_{\varepsilon})$; inference must use the cluster-robust sandwich of Section 3 and Super Learner cross-validation at the cluster level (Supplement, Section S3).
 2.1 Potential Outcomes and the Causal Target
-For intervention value $a\in\{0,1\}$, let $S_{ijt}(a)$ and $Y_{ijt}(a)$ denote the potential outcomes obtained by the intervention $\mathrm{do}(A_{jt}=a)$ in the SCM.
-The causal Average Treatment Effect (ATE) is:
-$\Psi(P_{0})\;=\;\mathbb{E}\bigl[Y_{ijt}(1)-Y_{ijt}(0)\bigr],$ (7)
-where the expectation is over the marginal distribution of $(W_{ijt},b_{j},t)$ in the trial. The identification of $\Psi(P_{0})$ from $P_{0}$ is the subject of Section 2.2.
+For $a\in\{0,1\}$, the causal ATE is $\Psi(P_{0})=\mathbb{E}[Y_{ijt}(1)-Y_{ijt}(0)]$, where the expectation is over the marginal distribution of $(W_{ijt},b_{j},t)$.
 2.2 Non-Parametric Identification
-The central identification insight is that the target causal effect can be represented through a surrogate bridge: one first learns the observed-outcome regression among units with realized primary outcomes and then integrates that regression over the treatment-specific surrogate distribution.
-This representation identifies the causal estimand without placing inverse observation weights in the target functional itself, which is precisely what makes it attractive in regimes where observation probabilities become small.
-We establish the following assumptions to formalize this representation.
 Assumption 1 (Consistency).
-For any observed treatment assignment $A_{jt}=a$,  $S_{ijt}=S_{ijt}(a)$ and $Y_{ijt}=Y_{ijt}(a)$.
+For $A_{jt}=a$: $S_{ijt}=S_{ijt}(a)$ and $Y_{ijt}=Y_{ijt}(a)$.
 Assumption 2 (Sequential Randomization).
-Conditional on the cluster history, treatment is independent of potential outcomes:
-$\bigl(Y(0),Y(1),S(0),S(1)\bigr)\;\perp\;A_{jt}\;\big|\;W_{ijt},\,t.$ (8)
-This holds by design: $A_{jt}=I(t\geq\tau_{j})$ and $\tau_{j}$ is randomized independently of individual outcomes.
-Assumption 3 (Surrogate-Mediated Missing at Random).
-The unobserved outcome provides no additional information about the censoring probability once the surrogate is known (Rubin, 1976):
-$P(\Delta_{ijt}=1\mid Y_{ijt},S_{ijt},A_{jt},W_{ijt},t)\;=\;P(\Delta_{ijt}=1\mid S_{ijt},A_{jt},W_{ijt},t).$ (9)
-This is structurally encoded in the SCM by the absence of the edge $Y_{ijt}\to\Delta_{ijt}$ (Figure 1, Equation (5)).
-Remark 2 (Practical plausibility of Assumption 3).
-Assumption 3 is most defensible when missingness is administrative (calendar-driven) rather than informative (health-driven).
-In the SW-CRT setting, the dominant censoring mechanism — the trial’s fixed end date — is independent of $Y_{ijt}$ conditional on $S_{ijt}$.
-A strong surrogate candidate satisfies three properties: (i) causal proximity: $S$ lies on or near the pathway $A\to S\to Y$; (ii) temporal availability: $S$ is observed for all individuals regardless of crossover time; and (iii) biological stationarity: $\mathbb{E}[Y\mid S,A,W,t]$ does not change systematically over calendar time.
-Examples include 3-month viral load predicting 12-month suppression in HIV trials, and discharge-process scores predicting 30-day readmission in quality-improvement SW-CRTs.
-Assumption 4 (Support Positivity for the Observed Outcome Regression).
-There exists a constant $c>0$ such that, for all $(s,a,w,t)$ in the support of $(S,A,W,t)$ relevant to the target parameter,
-$P(\Delta_{ijt}=1\mid S_{ijt}=s,A_{jt}=a,W_{ijt}=w,t)\;\geq\;c.$ (10)
-Remark 3 (Support positivity versus inverse-weighting regularity).
-Assumption 4 is a support condition for identifying $E[Y\mid S,A,W,t]$ from the observed cohort $\{\Delta=1\}$, not an inverse-weighting regularity condition.
-IPCW-type estimators are sensitive to small values of $g_{\Delta}$ because $g_{\Delta}^{-1}$ appears directly in the influence function, so that even when Assumption 4 holds with a small constant $c$, the IPCW estimating equation becomes highly variable.
-The surrogate-bridge parameter, by contrast, does not contain $g_{\Delta}^{-1}$ in the target functional; it requires only that the complete-case outcome regression $E[Y\mid S,A,W,t,\Delta=1]$ be well-defined on the support over which the surrogate-bridge formula integrates.
+$(Y(0),Y(1),S(0),S(1))\perp A_{jt}\mid W_{ijt},t$. This holds by design since $\tau_{j}$ is randomized independently of outcomes.
+Assumption 3 (Surrogate-Mediated MAR).
+$P(\Delta_{ijt}=1\mid Y_{ijt},S_{ijt},A_{jt},W_{ijt},t)=P(\Delta_{ijt}=1\mid S_{ijt},A_{jt},W_{ijt},t)$ (Rubin, 1976).
+Assumption 3 is most defensible when censoring is calendar-driven rather than health-driven; strong surrogates satisfy causal proximity ($S$ on the path $A\to S\to Y$), temporal availability, and biological stationarity of $\mathbb{E}[Y\mid S,A,W,t]$.
+Assumption 4 (Support Positivity).
+$P(\Delta_{ijt}=1\mid S_{ijt}=s,A_{jt}=a,W_{ijt}=w,t)\geq c>0$ on the relevant support.
+Unlike IPCW, the surrogate-bridge target functional does not contain $g_{\Delta}^{-1}$; this is a support condition for identifying $\mathbb{E}[Y\mid S,A,W,t]$ from $\{\Delta=1\}$, not an inverse-weighting regularity condition.
 Assumption 5 (Design-Based Treatment Support).
-The marginal treatment probability, averaged over the trial design space, is non-degenerate:
-$0\;<\;\mathbb{E}_{t}\bigl[g_{A}(1\mid W,t)\bigr]\;<\;1.$ (11)
-Remark 4 (Why Treatment Positivity Holds Marginally Despite Pointwise Violations).
-At any fixed time $t$ and cluster $j$, the treatment assignment is deterministic: $A_{jt}=I(t\geq\tau_{j})$, so pointwise positivity $0<P(A_{jt}=1\mid W,t)<1$ fails everywhere.
-A naive reading would therefore conclude that treatment positivity is structurally violated by the SW-CRT design.
-This is resolved by observing that the identification formula (1) marginalizes over the trial’s design space, averaging jointly over both $W$ and $t$.
-The design-based propensity is known by construction from the randomization schedule.
-Under a uniform randomization of crossover times:
-$g_{A}(1\mid W,t)\;=\;P(\tau_{j}\leq t)\;=\;\frac{t-1}{T-1}.$ (12)
-Averaging over the interior steps $t\in\{2,\ldots,T-1\}$ yields $\mathbb{E}_{t}[g_{A}(1\mid W,t)]=1/2$, so Assumption 5 holds.
-In practice, boundary steps $t=1$ and $t=T$ should be excluded, as $g_{A}$ equals 0 and 1 respectively at these endpoints.
-A further consequence is that $g_{A}$ need not be estimated from the data.
-In Section 4, we plug in the design-based propensity directly, eliminating one source of nuisance estimation error from the second-order remainder $R_{2}$ (Supplement, Section S2).
+$0<\mathbb{E}_{t}[g_{A}(1\mid W,t)]<1$. Under uniform crossover randomization $g_{A}(1\mid W,t)=(t-1)/(T-1)$, known by construction, so $g_{A}$ is plugged in directly rather than estimated.
 Theorem 1 (Identification via a Surrogate Bridge under Support Positivity).
@@
 Proof.
-By Assumptions 1 (Consistency) and 2 (Sequential Randomization), the potential outcome mean satisfies:
-$\mathbb{E}[Y(a)\mid S(a),W,t]\;=\;\mathbb{E}[Y\mid S,A=a,W,t].$ (14)
-By Assumption 3 (Surrogate-Mediated MAR) and Assumption 4 (Support Positivity), we may condition on the observed cohort ($\Delta=1$):
-$\mathbb{E}[Y\mid S,A=a,W,t]\;=\;\mathbb{E}[Y\mid S,A=a,W,t,\Delta=1]\;\equiv\;\bar{Q}_{Y}(S,a,W,t).$ (15)
-Integrating $\bar{Q}_{Y}$ over the conditional surrogate distribution:
-$\mathbb{E}[Y(a)\mid W=w,t]\;=\;\int\bar{Q}_{Y}(s,a,w,t)\,dP(s\mid a,w,t)\;\equiv\;\bar{Q}_{\mathrm{int}}(a,w,t).$ (16)
-Marginalizing over the trial design space yields (1).∎∎
+By Assumptions 12, $\mathbb{E}[Y(a)\mid S(a),W,t]=\mathbb{E}[Y\mid S,A=a,W,t]$.
+By Assumption 3 and 4, conditioning on $\{\Delta=1\}$ is valid: $\mathbb{E}[Y\mid S,A=a,W,t]=\mathbb{E}[Y\mid S,A=a,W,t,\Delta=1]\equiv\bar{Q}_{Y}(S,a,W,t)$.
+Integrating over $P(S\mid a,W,t)$ and marginalizing over $(W,t)$ yields (1).∎∎
 3 Semiparametric Theory
@@
 $g_{\Delta}(1\mid S,A,W,t)$ $=P(\Delta=1\mid S,A,W,t).$ (26)
-Remark 6 (Observed-Data Regression vs. Counterfactual Evaluation of $\bar{Q}_{\mathrm{int}}$).
-Equation (31) defines $\bar{Q}_{\mathrm{int}}(A,W,t)$ as an observed-data regression of $\bar{Q}_{Y}$ on the observed treatment $A$.
-During estimation, this regression is fit on the full data using the observed $A\in\{0,1\}$.
-Counterfactual predictions are then obtained by evaluating the fitted model at the fixed intervention values $A=1$ and $A=0$ separately, yielding $\hat{\bar{Q}}_{\mathrm{int}}(1,W,t)$ and $\hat{\bar{Q}}_{\mathrm{int}}(0,W,t)$.
-These counterfactual evaluations enter the ATE estimator (37).
-The regression is fit on the observed scale and evaluated on the interventional scale; conflating these two steps would produce a biased plug-in estimator, which is precisely what the TMLE fluctuation step in Section 4.2 corrects.
-As established in Remark 4, the treatment propensity $g_{A}(1\mid W,t)=(t-1)/(T-1)$ is known from the randomization schedule and is plugged in directly.
-Only $g_{\Delta}$ requires data-adaptive estimation.
+$\bar{Q}_{\mathrm{int}}$ is fit on the observed $A$ and evaluated counterfactually at $A=1$ and $A=0$ to produce $\hat{\bar{Q}}_{\mathrm{int}}(1,W,t)$ and $\hat{\bar{Q}}_{\mathrm{int}}(0,W,t)$; the TMLE fluctuation step corrects the resulting plug-in bias.
+The treatment propensity $g_{A}(1\mid W,t)=(t-1)/(T-1)$ is plugged in directly from the randomization schedule; only $g_{\Delta}$ requires data-adaptive estimation.
 4.2 Stage 2: The Nested Fluctuation Step
@@
 Proof.
-Asymptotic linearity (43) follows from the von Mises expansion of $\hat{\Psi}_{\mathrm{TMLE}}$ around $P_{0}$.
-The TMLE fluctuation step is constructed precisely to solve the efficient score equation $J^{-1}\sum_{j}\mathrm{EIC}_{j}=0$ (up to the convergence tolerance (36)), which eliminates the first-order bias term.
-The residual is the second-order remainder $R_{2}(\hat{P},P_{0})$ characterized in Supplement, Section S2.4.
-Under condition (C3), $R_{2}=o_{P}(J^{-1/2})$; this follows from the bilinear form of $R_{2}$ established in Supplement, Section S2.4 and the Cauchy–Schwarz inequality applied to each product of $L^{2}(P_{0})$ errors.
-When the CV-TMLE of Supplement, Section S3 is used, condition (C3) is achieved without requiring a Donsker condition on the nuisance function classes, because sample splitting eliminates the empirical process terms that would otherwise arise (Zheng and van der Laan, 2011).
-Asymptotic normality (44) and consistency of the sandwich variance estimator then follow from the Lindeberg–Feller CLT applied to the i.i.d.
-cluster scores $\{\mathrm{EIC}_{j}\}_{j=1}^{J}$; the complete argument is given in Supplement, Section S2.4.∎∎
-Remark 7 (Double robustness of the asymptotic result).
-The product-rate condition (C3) is doubly robust in the following sense (Bang and Robins, 2005; Tsiatis, 2006): it is satisfied if either (i) the outcome models $(\bar{Q}_{Y},\bar{Q}_{\mathrm{int}})$ are consistently estimated at any rate, or (ii) the propensity mechanisms $(g_{A},g_{\Delta})$ are consistently estimated at any rate.
-In particular, since $g_{A}$ is known by design in the SW-CRT (Remark 4), the condition reduces to requiring only that $\hat{\bar{Q}_{Y}}$ and $\hat{\bar{Q}_{\mathrm{int}}}$ converge at rate $o_{P}(J^{-1/4})$, a strictly weaker requirement than pointwise consistency.
-Proposition 2 (Nested Cross-Product Term in the DML Remainder).
-Let $f_{S}(\cdot\mid a,W,t)\equiv P_{0}(S\in\cdot\mid A=a,W,t)$ denote the conditional density of the surrogate, and let $\hat{f}_{S}$ be any estimator of $f_{S}$ independent of the SA-TMLE nuisance estimators.
-Define the plug-in integral estimator $\hat{\bar{Q}}^{\,\mathrm{DML}}_{\mathrm{int}}(a,W,t)\equiv\int\hat{\bar{Q}}_{Y}(s,a,W,t)\,\hat{f}_{S}(s\mid a,W,t)\,d\mu(s)$, which constructs $\bar{Q}_{\mathrm{int}}$ as an integral of $\hat{\bar{Q}}_{Y}$ against $\hat{f}_{S}$.
-For the DML one-step estimator $\hat{\Psi}^{\mathrm{DML}}=\hat{\Psi}^{\mathrm{DML}}_{\mathrm{plug\text{-}in}}+N^{-1}\sum_{j,i,t}D^{*}(O_{ijt};\hat{\bar{Q}}_{Y},\hat{\bar{Q}}^{\,\mathrm{DML}}_{\mathrm{int}},\hat{g}_{A},\hat{g}_{\Delta})$ based on this construction, the von Mises expansion yields:
-$\hat{\Psi}^{\mathrm{DML}}-\Psi(P_{0})\;=\;\frac{1}{J}\sum_{j=1}^{J}\mathrm{EIC}_{j}\;+\;R_{2}^{\mathrm{TMLE}}(\hat{P},P_{0})\;+\;R_{SY}(\hat{P},P_{0})\;+\;o_{P}(J^{-1/2}),$ (45)
-where $R_{2}^{\mathrm{TMLE}}$ is the TMLE remainder (Supplement, Section S2.4) and the nested cross-product remainder is
-$R_{SY}(\hat{P},P_{0})\;=\;\sum_{a\in\{0,1\}}(-1)^{1-a}\iint\bigl(\hat{\bar{Q}}_{Y}-\bar{Q}^{0}_{Y}\bigr)(s,a,w,t)\,\bigl(\hat{f}_{S}-f^{0}_{S}\bigr)(s\mid a,w,t)\;d\mu(s)\,dP_{0}(w,t).$ (46)
-For $R_{SY}=o_{P}(J^{-1/2})$ the DML estimator requires the additional product-rate condition
-$\sum_{a\in\{0,1\}}\bigl\|\hat{\bar{Q}}_{Y}(\cdot,a,\cdot,\cdot)-\bar{Q}^{0}_{Y}(\cdot,a,\cdot,\cdot)\bigr\|_{P_{0}}\cdot\bigl\|\hat{f}_{S}(\cdot\mid a,\cdot,\cdot)-f^{0}_{S}(\cdot\mid a,\cdot,\cdot)\bigr\|_{P_{0}}\;=\;o_{P}(J^{-1/2}),$ (47)
-which involves estimating the conditional density $f_{S}$ at rate $o_{P}(J^{-1/4})$ in $L^{2}(P_{0})$.
-The SA-TMLE eliminates $R_{SY}$ without estimating $f_{S}$: the second fluctuation step (Section 4.2) enforces $N^{-1}\sum_{j,i,t}D_{S,a}(O_{ijt})=0$, which constrains $\hat{\bar{Q}}^{\,*}_{\mathrm{int}}$ to the propensity-weighted empirical conditional mean of $\hat{\bar{Q}}^{\,*}_{Y}$, thereby absorbing $R_{SY}$ into the efficient score without requiring knowledge of $f_{S}$.
-The proof is given in Supplement, Section S2.6.
-Remark 8 (Why DML cross-fitting does not eliminate $R_{SY}$).
-[…]
+The von Mises expansion of $\hat{\Psi}_{\mathrm{TMLE}}$ around $P_{0}$ yields
+$\hat{\Psi}_{\mathrm{TMLE}}-\Psi(P_{0})\;=\;\frac{1}{J}\sum_{j=1}^{J}\mathrm{EIC}_{j}\;+\;R_{2}(\hat{P},P_{0})\;+\;R_{SY}(\hat{P},P_{0})\;+\;o_{P}(J^{-1/2}),$ (38)
+where $R_{2}$ is the standard bilinear nuisance-error remainder and $R_{SY}$ is the nested cross-product term of Proposition 1.
+Elimination of $R_{SY}$.
+The second fluctuation step (Section 4.2) enforces the score equation $N^{-1}\sum_{j,i,t}D_{S,a}(O_{ijt})=0$ for each $a$.
+Expanding $D_{S,a}$ from (11) shows that this constraint sets the propensity-weighted empirical mean of $\bar{Q}^{\,*}_{Y}$ equal to $\bar{Q}^{\,*}_{\mathrm{int}}$ at the observed surrogate values, i.e.,
+$\frac{1}{N}\sum_{j,i,t}\frac{I(A=a)}{\hat{g}_{A}(a\mid W,t)}\bigl[\bar{Q}^{\,*}_{Y}(S_{ijt},a,W_{ijt},t)-\bar{Q}^{\,*}_{\mathrm{int}}(a,W_{ijt},t)\bigr]\;=\;0.$ (39)
+Because $\bar{Q}^{\,*}_{\mathrm{int}}$ is thereby defined as a propensity-weighted empirical average of $\bar{Q}^{\,*}_{Y}$ over the observed surrogate distribution—rather than as an explicit integral against an estimated $\hat{f}_{S}$—the cross-product $(\hat{Q}_{Y}-Q^{0}_{Y})(\hat{f}_{S}-f^{0}_{S})$ that constitutes $R_{SY}$ never enters the expansion; $R_{SY}$ is absorbed into the efficient score without any requirement on the rate of $\hat{f}_{S}$.
+The formal calculation is in Supplement, Section S2.6.
+Control of $R_{2}$.
+The residual remainder $R_{2}(\hat{P},P_{0})$ is bilinear in the nuisance errors $(\hat{Q}_{Y}-Q^{0}_{Y})$ and $(\hat{g}_{\Delta}-g^{0}_{\Delta})$, and in $(\hat{Q}_{\mathrm{int}}-Q^{0}_{\mathrm{int}})$ and $(\hat{g}_{A}-g^{0}_{A})$.
+Cauchy–Schwarz gives $|R_{2}|\leq\|\hat{Q}_{Y}-Q^{0}_{Y}\|_{P_{0}}\|\hat{g}_{\Delta}-g^{0}_{\Delta}\|_{P_{0}}+\|\hat{Q}_{\mathrm{int}}-Q^{0}_{\mathrm{int}}\|_{P_{0}}\|\hat{g}_{A}-g^{0}_{A}\|_{P_{0}}$, which is $o_{P}(J^{-1/2})$ by condition (C3).
+When the CV-TMLE of Supplement, Section S3 is used, sample splitting eliminates the empirical process terms that would otherwise require a Donsker condition (Zheng and van der Laan, 2011).
+Asymptotic normality.
+With $R_{2}+R_{SY}=o_{P}(J^{-1/2})$ established, the Lindeberg–Feller CLT applied to the i.i.d.
+cluster scores $\{\mathrm{EIC}_{j}\}_{j=1}^{J}$ gives (37).
+Consistency of the sandwich estimator (32) for $\sigma^{2}/J$ follows by the law of large numbers applied at the cluster level.
+The complete argument is in Supplement, Section S2.4.∎∎
+Remark 3 (Double robustness and role of $R_{SY}$).
+Condition (C3) is doubly robust (Bang and Robins, 2005): it holds if either the outcome models $(\bar{Q}_{Y},\bar{Q}_{\mathrm{int}})$ or the propensities $(g_{A},g_{\Delta})$ are consistently estimated.
+Since $g_{A}$ is known by design, (C3) reduces to $\|\hat{\bar{Q}}_{Y}-\bar{Q}^{0}_{Y}\|_{P_{0}}\cdot\|\hat{g}_{\Delta}-g^{0}_{\Delta}\|_{P_{0}}=o_{P}(J^{-1/2})$.
+The key separation from DML is (39): the second fluctuation constrains $\bar{Q}^{\,*}_{\mathrm{int}}$ to a propensity-weighted mean of $\bar{Q}^{\,*}_{Y}$ over observed surrogates, so the cross-product $(\hat{Q}_{Y}-Q^{0}_{Y})(\hat{f}_{S}-f^{0}_{S})$ constituting $R_{SY}$ never enters the expansion.
 5.3 Non-asymptotic Coverage Bound
-Beyond asymptotic validity, the bridge-based estimator admits an explicit non-asymptotic coverage analysis at the cluster level.
-The resulting Berry–Esseen bound (Theorem S1, Supplementary Material) quantifies how bounded cluster scores, variance concentration, and nuisance-estimation remainder jointly determine finite-sample undercoverage, thereby linking the semiparametric theory directly to the small-$J$ regimes encountered in practice.
-Corollary S1 gives the minimum cluster count $J^{*}$ for a target coverage level; for the simulation DGP of Section 6, $J^{*}\approx 27$ for 1% maximum undercoverage.
-Corollary S2 (Supplement, Section S2.7) extends this to characterize how nuisance convergence rates govern the sandwich variance gap, showing that the constant coverage shortfall observed in Block I simulations is the finite-sample signature of estimation near the product-rate boundary.
+A Berry–Esseen bound (Theorem S1, Supplement) quantifies finite-sample undercoverage of the sandwich interval in terms of bounded cluster scores, variance concentration, and the nuisance remainder, giving a minimum cluster count $J^{*}\approx 27$ for 1% maximum undercoverage at the simulation DGP.
+Corollary S2 shows that the constant coverage shortfall in Block I is the finite-sample signature of estimation near the product-rate boundary: the fluctuation-step variance contributes an $O(J^{-1})$ term not captured by the sandwich, and this term does not shrink relative to the leading variance term under parametric nuisance rates.
+The leave-one-cluster-out jackknife for the full SA-TMLE estimator resolves this gap in practice (Section 6.3), yielding near-nominal coverage at $J\geq 30$.
+Formal jackknife variance consistency for targeted estimators with data-adaptive nuisance follows from the arguments of Zheng and van der Laan (2011); a complete non-asymptotic treatment is deferred to future work.
 6 Monte Carlo Simulation Study
-The simulation study is designed to examine the main qualitative implications of the theory: relative stability of the surrogate-bridge estimator as observation probabilities become small (Blocks I and III), the finite-sample behavior of the asymptotic double-robustness theory under deliberate nuisance misspecification (Block II), and the relationship between coverage and the remainder-variance analysis of Section 5.3.
-The three scenario blocks are organized around increasing administrative censoring severity, misspecified nuisance models, and varying cluster counts.
-Throughout, “double robustness” refers to the asymptotic identification of the first-order remainder; finite-sample performance can still depend strongly on which nuisance component is misspecified and on the magnitude of the residual second-order error.
+Three simulation blocks examine the main implications of the theory: stability under increasing censoring (Blocks I, III), finite-sample double-robustness under nuisance misspecification (Block II), and coverage scaling with $J$ (Block I).
+Throughout, “double robustness” refers to asymptotic first-order identification; finite-sample performance depends on which nuisance component is misspecified and the magnitude of the residual second-order error.
 6.1 Data-Generating Process
-Trial design.
-We simulate a standard SW-CRT with $J$ clusters, $T=7$ calendar time steps, and $n_{j}=40$ individuals per cluster per step, giving a total sample size of $N=J\times T\times n_{j}$.
-Crossover times $\tau_{j}$ are drawn uniformly without replacement from $\{2,3,4,5,6,7\}$, so that exactly J/(T1)J/(T-1) clusters (rounded to the nearest integer) cross over at each step.
-All clusters begin in the control condition: $A_{jt}=\mathbf{1}(t\geq\tau_{j})$.
-Administrative censoring of the delayed outcome is structurally imposed: an individual in cluster $j$ at time step $t$ can contribute an observed delayed outcome only if $\tau_{j}+t_{\mathrm{lag}}\leq T+t_{\mathrm{grace}}$, where $t_{\mathrm{lag}}=2$ steps is the outcome delay and $t_{\mathrm{grace}}=1$ step is an administrative grace period.
-This implies that clusters crossing over at steps T1T-1 and $T$ have near-zero probability of observing $Y$, creating the design-induced near-boundary observation regime described in Section 1.
-Cluster random effect.
-The cluster-level random effect is drawn independently as $b_{j}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\sigma_{b}^{2})$ with $\sigma_{b}=0.184$ ($\sigma_{b}^{2}=0.034$, $\mathrm{ICC}\approx 0.05$), consistent with the ICC model of Section 2.3 and empirical ICC levels typical of public-health SW-CRTs (Hemming and Taljaard, 2020).
-Baseline covariates.
-Each individual has a $p=3$ dimensional baseline covariate vector:
-$W_{ij}=(W_{ij,1},\,W_{ij,2},\,W_{ij,3})$
-where $W_{ij,1}\sim\mathcal{N}(0,1)$, $W_{ij,2}\sim\mathrm{Bernoulli}(0.4)$, and $W_{ij,3}\sim\mathrm{Uniform}(0,1)$, all independent across individuals and clusters.
-Secular time trend.
-We specify the true secular time trend as a non-linear function of the calendar step to represent the realistic scenario where a linear time trend is misspecified:
-$\lambda(t)=0.5\sin\!\left(\frac{\pi(t-1)}{T-1}\right)+0.3\left(\frac{t-1}{T-1}\right)^{2}.$
-This ensures that parametric estimators using a simple linear-in-$t$ fixed effect will exhibit systematic secular time confounding bias.
-Short-term surrogate.
-The surrogate is generated from a linear model with a cluster random effect, treatment effect, covariate effects, and secular time trend:
-$S_{ijt}$ $=0.8\,A_{jt}+0.4\,W_{ij,1}-0.3\,W_{ij,2}+0.5\,\lambda(t)+0.6\,b_{j}+\varepsilon^{S}_{ijt},$
-$\varepsilon^{S}_{ijt}$ $\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\,0.5^{2}).$ (48)
-The coefficient $0.6$ on $b_{j}$ induces a strong within-cluster surrogate–outcome correlation, making $S$ a useful bridge for the missing delayed outcome.
-Observation indicator.
-Administrative censoring follows a logistic model consistent with Assumption 3:
-$\operatorname{logit}\bigl(\mathbb{P}(\Delta_{ijt}=1\mid S_{ijt},A_{jt},W_{ij},t)\bigr)$ $=3.0-2.5\cdot\mathbf{1}(t\geq T-1)-0.8\cdot\mathbf{1}(t=T)$
-$\quad+0.4\,S_{ijt}+0.3\,A_{jt}-0.2\,W_{ij,3}.$ (49)
-The large negative late-step coefficients concentrate censoring in the final two steps (overall rate $\approx 28\%$, rising to $\approx 80\%$ at $t=T$), creating the near-boundary observation regime for late-crossing clusters.
-Delayed primary outcome.
-Following the structural equation (6) and the ICC decomposition of Section 2.3:
-$Y_{ijt}$ $=\Psi_{0}\,A_{jt}+0.5\,S_{ijt}+0.4\,W_{ij,1}-0.2\,W_{ij,2}+0.6\,W_{ij,3}+0.8\,\lambda(t)+b_{j}+\varepsilon^{Y}_{ijt},$
-$\varepsilon^{Y}_{ijt}$ $\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\,0.8^{2}),$ (50)
-where $\Psi_{0}=-0.28$ is the direct-effect coefficient, $\mathrm{ICC}=0.034/(0.034+0.64)\approx 0.05$.
-True ATE.
-The true total causal ATE is $\Psi^{*}=-0.28+0.5\times 0.8=0.12$ (direct effect plus mediated path through $S$), placed near the detection boundary for $J\approx 30$.
+We simulate a SW-CRT with $J$ clusters, $T=7$ calendar time steps, and $n_{j}=40$ individuals per cluster per step ($N=J\times T\times n_{j}$).
+Crossover times $\tau_{j}$ are drawn uniformly from $\{2,\ldots,7\}$; administrative censoring is imposed structurally so that clusters crossing at steps T1T-1 and $T$ have near-zero $g_{\Delta}$ (outcome delay $t_{\mathrm{lag}}=2$, grace period $t_{\mathrm{grace}}=1$, overall censoring rate $\approx 28\%$).
+The cluster random effect satisfies $b_{j}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,0.034)$ ($\mathrm{ICC}\approx 0.05$).
+Baseline covariates are $W_{ij,1}\sim\mathcal{N}(0,1)$, $W_{ij,2}\sim\mathrm{Bernoulli}(0.4)$, $W_{ij,3}\sim\mathrm{Uniform}(0,1)$.
+The secular trend is non-linear, $\lambda(t)=0.5\sin(\pi(t-1)/(T-1))+0.3((t-1)/(T-1))^{2}$, so that parametric linear-in-$t$ models are misspecified.
+The surrogate satisfies $S_{ijt}=0.8A_{jt}+0.4W_{ij,1}-0.3W_{ij,2}+0.5\lambda(t)+0.6b_{j}+\varepsilon^{S}_{ijt}$, $\varepsilon^{S}_{ijt}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,0.25)$; the coefficient $0.6$ on $b_{j}$ ensures $S$ is an informative bridge.
+The observation indicator and outcome follow
+$\operatorname{logit}P(\Delta_{ijt}=1\mid S,A,W,t)$ $=3.0-2.5\cdot\mathbf{1}(t\geq T-1)-0.8\cdot\mathbf{1}(t=T)+0.4S_{ijt}+0.3A_{jt}-0.2W_{ij,3},$ (40)
+$Y_{ijt}$ $=\Psi_{0}A_{jt}+0.5S_{ijt}+0.4W_{ij,1}-0.2W_{ij,2}+0.6W_{ij,3}+0.8\lambda(t)+b_{j}+\varepsilon^{Y}_{ijt},$ (41)
+with $\varepsilon^{Y}_{ijt}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,0.64)$ and direct effect $\Psi_{0}=-0.28$.
+The true ATE is $\Psi^{*}=-0.28+0.5\times 0.8=0.12$, placed near the detection boundary for $J\approx 30$.
 6.2 Simulation Scenarios
@@
 Block II: Misspecified nuisance models.
-$J=30$, $\Psi^{*}=0.12$. We consider three misspecification scenarios designed to probe the double robustness property of Theorem 2:
-  1. (i)
-    Outcome model misspecified, propensities correctly specified.
-The outcome models $\bar{Q}_{Y}$ and $\bar{Q}_{\mathrm{int}}$ are fit using a main-effects linear regression that omits the non-linear time trend $\lambda(t)$ and the interaction $S_{ijt}\times A_{jt}$; $g_{\Delta}$ is estimated by the correctly specified logistic model.
-Under the asymptotic double-robustness theory, consistency is still possible when the propensity side is correctly specified, but finite-sample performance may deteriorate if the misspecified outcome model leaves a large structured remainder.
-  2. (ii)
-    Propensities misspecified, outcome models correctly specified.
-$g_{\Delta}$ is estimated by logistic regression omitting $S_{ijt}$ (violating the surrogate-conditional independence structure of Assumption 3); $\bar{Q}_{Y}$ and $\bar{Q}_{\mathrm{int}}$ are correctly specified.
-Because $g_{A}$ is design-based and used as a plug-in (Remark 3), the treatment propensity remains correctly specified; only $g_{\Delta}$ is misspecified.
-Under the asymptotic double-robustness theory, the estimator is expected to remain stable when the outcome regressions are correctly specified, even if $g_{\Delta}$ is misspecified.
-  3. (iii)
-    Both nuisance models misspecified.
-Both the outcome models and $g_{\Delta}$ are misspecified as in (i) and (ii).
-This scenario is expected to produce estimator bias and is included as a negative control to verify that double robustness does not extend to simultaneous misspecification of both sides.
+$J=30$.
+Three scenarios probe Theorem 2’s double-robustness: (i) outcome models misspecified (non-linear trend and $S\times A$ interaction omitted), $g_{\Delta}$ correct; (ii) $g_{\Delta}$ misspecified (surrogate $S$ omitted from the censoring model), outcome models correct; (iii) both misspecified.
+Scenarios (i) and (iii) are expected to show bias; scenario (ii) tests the arm of double robustness where consistent outcome models compensate for $g_{\Delta}$ misspecification.
 Block III: Increasing administrative censoring severity.
@@
 Block I: Baseline performance across cluster counts.
-Table 3 reports performance metrics for all three estimators across $J\in\{10,20,30,50,100\}$.
-The CV-TMLE achieves near-zero bias (<0.004<0.004) at all cluster counts.
-The GLMM exhibits persistent bias of approximately +0.07+0.07 at $J=30$ due to the misspecified secular time trend, and its coverage collapses from 0.52 at $J=10$ to 0.07 at $J=100$ as the standard error shrinks around the biased point estimate.
-IPCW exhibits substantial positive bias (approximately +0.22+0.22 at $J=30$) and markedly higher variance than the CV-TMLE, consistent with the theoretical prediction that inverse weighting becomes unstable when $g_{\Delta}$ concentrates near the positivity boundary for late-crossing clusters.
-Coverage for CV-TMLE stabilises near 0.87–0.91 across $J\geq 20$, with no systematic improvement as $J$ grows to 100.
-Table 5 decomposes this gap: the sandwich estimates $\mathrm{Var}(\mathrm{EIC}_{j})/J$ but does not capture the variance of the second-order remainder $R_{2}$.
-The empirical variance exceeds the mean sandwich by a factor of about 1.6 at all $J$; this magnitude is broadly consistent with the observed coverage shortfall in the 0.87–0.89 range (see predicted column).
-This ratio is approximately constant because both terms are $O(J^{-1})$ under parametric nuisance rates.
-The theory suggests that this gap should diminish once nuisance estimation is sufficiently faster than the product-rate boundary; the present simulations indicate that the parametric nuisance specifications used here remain in a regime where that improvement is not yet visible.
-CV-TMLE power scales from 0.41 at $J=10$ to 1.00 at $J=100$, reflecting strong signal recovery together with very small empirical bias.
-Table 2: Variance decomposition for CV-TMLE coverage (Block I).
-$\widehat{\mathrm{Var}}_{\mathrm{emp}}$: empirical variance of $\hat{\Psi}$ across 1,000 replicates.
+Table S5 (Supplement) reports point-estimation metrics for all three estimators across $J\in\{10,20,30,50,100\}$; the inference comparison between sandwich and jackknife for CV-TMLE is in Table 2.
+Point estimation.
+CV-TMLE achieves near-zero bias ($|\mathrm{Bias}|<0.004$) at all cluster counts, with RMSE improving steadily from 0.092 at $J=10$ to 0.030 at $J=100$.
+GLMM exhibits persistent bias of approximately +0.07+0.07 due to the misspecified secular time trend; its coverage collapses from 0.52 at $J=10$ to near zero at $J=100$ as the standard error shrinks around the biased point estimate.
+IPCW exhibits substantial positive bias (approximately +0.22+0.22) and markedly higher variance, consistent with the theoretical prediction that inverse weighting becomes unstable when $g_{\Delta}$ concentrates near zero for late-crossing clusters.
+Sandwich variance and undercoverage.
+CV-TMLE with the sandwich interval achieves coverage of 0.87–0.89 across all $J$, well below the nominal 0.95.
+Table 1 diagnoses the source: the sandwich estimates $\mathrm{Var}(\mathrm{EIC}_{j})/J$ but omits the variance contribution from the fluctuation-step parameters $(\varepsilon_{Y},\varepsilon_{S})$, which vary across leave-one-cluster-out fits.
+The empirical variance exceeds the mean sandwich by a factor of approximately 1.6 at all $J$; this ratio is stable because both the sandwich and the omitted term are $O(J^{-1})$ under the present nuisance specification, so neither diminishes faster than the other.
+Jackknife variance and coverage restoration.
+The leave-one-cluster-out jackknife, applied to the full two-stage estimator (re-running both fluctuation steps for each held-out cluster), captures the fluctuation-step variance that the sandwich misses.
+Table 1 shows that the mean jackknife variance tracks the empirical variance closely at all $J$.
+The resulting coverage (Table 2, Jackknife columns) is near-nominal at the practically important cluster sizes: 0.956 at $J=30$ and 0.960 at $J=50$.
+At very small $J$ (10 and 20) the jackknife is mildly conservative (0.970–0.980), as the leave-one-out procedure loses a larger fraction of training data per fold and slightly overestimates variability; this is a well-understood property of the jackknife in small samples and represents a safe failure mode.
+By $J\geq 30$ the conservatism resolves and coverage is stable near 0.956–0.968 through $J=100$.
+The power tradeoff is modest: at $J=30$, jackknife power is 0.57 versus sandwich power of 0.73, reflecting wider intervals.
+By $J=50$, jackknife power reaches 0.786, and by $J=100$ it is 0.970 — essentially the same as the sandwich at those sizes.
+The undercoverage was a variance-calibration issue, not a failure of the point estimator; the cluster jackknife resolves it without altering the bias or RMSE.
+Table 1: Variance decomposition for CV-TMLE coverage (Block I, 500 replicates).
+$\widehat{\mathrm{Var}}_{\mathrm{emp}}$: empirical variance of $\hat{\Psi}$.
 $\overline{\mathrm{Sand.}}$: mean sandwich variance estimate.
-Ratio >1>1 indicates sandwich underestimation attributable to unaccounted $R_{2}$ variance.
-Predicted coverage computed as $P(|Z|<t_{J-1}/\sqrt{\mathrm{Ratio}})$; agreement with actual coverage is approximate.
-$J$ $\widehat{\mathrm{Var}}_{\mathrm{emp}}$ ($\times 10^{-3}$) $\overline{\mathrm{Sand.}}$ ($\times 10^{-3}$) Ratio Predicted Cov. Actual Cov.
-10 8.44 4.80 1.76 0.912 0.878
-20 4.83 2.83 1.71 0.891 0.869
-30 2.93 1.84 1.59 0.895 0.892
-50 1.80 1.11 1.62 0.886 0.866
-100 0.87 0.54 1.61 0.883 0.878
-Table 3: Block I simulation results: performance across cluster counts ($\Psi^{*}=0.12$, $T=7$, $n_{j}=40$, 1,000 replicates). “Cov.” denotes 95% CI coverage ($t_{J-1}$ critical values).
-$J$ Estimator Bias Var $(\times 10^{-2})$ RMSE Cov. Power
-10 GLMM +0.121+0.121 0.41 0.137 0.519 0.970
-IPCW +0.235+0.235 0.61 0.248 0.985 0.172
-CV-TMLE +0.004+0.004 0.84 0.092 0.878 0.410
-20 GLMM +0.104+0.104 0.21 0.114 0.373 0.998
-IPCW +0.231+0.231 0.33 0.238 0.934 0.649
-CV-TMLE +0.001+0.001 0.48 0.070 0.869 0.554
-30 GLMM +0.066+0.066 0.16 0.077 0.597 0.999
-IPCW +0.219+0.219 0.22 0.224 0.787 0.939
-CV-TMLE +0.000+0.000 0.29 0.054 0.892 0.727
-50 GLMM +0.083+0.083 0.08 0.088 0.203 1.000
-IPCW +0.224+0.224 0.12 0.227 0.220 1.000
-CV-TMLE +0.003+0.003 0.18 0.043 0.866 0.897
-100 GLMM +0.074+0.074 0.05 0.077 0.068 1.000
-IPCW +0.220+0.220 0.06 0.221 0.000 1.000
-CV-TMLE +0.000+0.000 0.09 0.030 0.878 0.997
+$\overline{\mathrm{Jack.}}$: mean cluster-jackknife variance estimate.
+Ratio $\widehat{\mathrm{Var}}_{\mathrm{emp}}/\overline{\mathrm{Sand.}}$; values >1>1 indicate sandwich underestimation attributable to unaccounted fluctuation-step variance.
+The jackknife variance tracks the empirical variance closely, explaining the near-nominal jackknife coverage in Table 2.
+$J$ $\widehat{\mathrm{Var}}_{\mathrm{emp}}$ ($\times 10^{-3}$) $\overline{\mathrm{Sand.}}$ ($\times 10^{-3}$) $\overline{\mathrm{Jack.}}$ ($\times 10^{-3}$) Ratio Sand. Cov. Jack. Cov.
+10 8.44 4.80 9.39 1.76 0.878 0.980
+20 4.83 2.83 5.37 1.71 0.869 0.970
+30 2.93 1.84 3.23 1.59 0.892 0.956
+50 1.80 1.11 2.01 1.62 0.866 0.960
+100 0.87 0.54 0.97 1.61 0.878 0.968
+Table 2: Block I CV-TMLE results: point estimation and inference comparison (500 replicates, $\Psi^{*}=0.12$, $T=7$, $n_{j}=40$).
+Bias and RMSE are identical across both variance estimators by construction.
+The jackknife is the leave-one-cluster-out procedure applied to the full two-stage estimator including both fluctuation steps.
+Full results for GLMM and IPCW comparators in Table S5 (Supplement).
+MC standard error for coverage $\approx 0.010$.
+Sandwich Jackknife
+$J$ Bias RMSE $\widehat{\mathrm{Var}}_{\mathrm{emp}}$ ($\times 10^{-3}$) Cov. Power Cov. Power
+10 +0.002+0.002 0.090 8.15 0.878 0.410 0.980 0.204
+20 0.001-0.001 0.068 4.66 0.869 0.554 0.970 0.332
+30 +0.003+0.003 0.054 2.91 0.892 0.727 0.956 0.570
+50 +0.003+0.003 0.043 1.80 0.866 0.897 0.960 0.786
+100 0.001-0.001 0.029 0.85 0.878 0.997 0.968 0.970
 Block II: Misspecified nuisance models.
@@
 ICC heterogeneity (Block IV).
-Additional simulations varying $\sigma_{b}^{2}\in\{0.006,0.034,0.071,0.160\}$ (ICC $\approx$ 0.01–0.20) confirm that the cluster-robust sandwich estimator maintains valid coverage across the empirically observed ICC range in implementation-science SW-CRTs; full results are in Supplement, Section S4.
-Table 5: Variance decomposition for CV-TMLE coverage (Block I).
-$\widehat{\mathrm{Var}}_{\mathrm{emp}}$: empirical variance of $\hat{\Psi}$ across 1,000 replicates.
-$\overline{\mathrm{Sand.}}$: mean sandwich variance estimate.
-Ratio >1>1 indicates sandwich underestimation attributable to unaccounted $R_{2}$ variance.
-Predicted coverage computed as $P(|Z|<t_{J-1}/\sqrt{\mathrm{Ratio}})$; agreement with actual coverage is approximate.
-$J$ $\widehat{\mathrm{Var}}_{\mathrm{emp}}$ ($\times 10^{-3}$) $\overline{\mathrm{Sand.}}$ ($\times 10^{-3}$) Ratio Predicted Cov. Actual Cov.
-10 8.44 4.80 1.76 0.912 0.878
-20 4.83 2.83 1.71 0.891 0.869
-30 2.93 1.84 1.59 0.895 0.892
-50 1.80 1.11 1.62 0.886 0.866
-100 0.87 0.54 1.61 0.883 0.878
+Additional simulations varying $\sigma_{b}^{2}\in\{0.006,0.034,0.071,0.160\}$ (ICC $\approx$ 0.01–0.20) confirm that the jackknife restores near-nominal coverage (0.946–0.968) across the full empirically observed ICC range in implementation-science SW-CRTs, extending the Block I finding to the ICC heterogeneity setting; full results are in Supplement, Section S4.
 7 Design-Calibrated Illustration: Washington State EPT Trial
@@
 This is administratively censored for late-crossing LHJs: wave-4 LHJs ($t_{j}=4$) have under 12 months of post-crossover follow-up, yielding an administrative censoring rate of 86% in that wave; wave-3 LHJs are partially censored (40%) from period 3 onward; overall 33.7% of records have missing 12-month outcomes.
-EPT uptake — whether the index patient accepted patient-delivered partner therapy within 30 days of STI diagnosis — is complete for all subjects regardless of wave and satisfies the three surrogate criteria of Remark 2: causal proximity to the intervention, temporal availability before the 12-month outcome, and biological stationarity.
+EPT uptake — whether the index patient accepted patient-delivered partner therapy within 30 days of STI diagnosis — is complete for all subjects regardless of wave and satisfies the three surrogate criteria of Assumption 3: causal proximity to the intervention, temporal availability before the 12-month outcome, and biological stationarity.
 We generate a replication dataset from the SCM of Section 2, with parameters fixed to match the published summary statistics of Golden et al.
@@
 8 Discussion
-We proposed a Surrogate-Assisted TMLE linking the longitudinal G-computation formula (1) to the cluster-level EIC of Section 3, providing an asymptotically doubly robust, non-parametric solution to the administrative censoring of late-crossing clusters in SW-CRTs.
-Standard IPCW fails in this setting because the censoring propensity $g_{\Delta}$ can become arbitrarily small for late-crossing clusters (Remark 3), producing highly unstable weights.
-The Surrogate-Assisted TMLE bypasses this instability by bridging missing outcomes through $S$, which is observed for all individuals regardless of the censoring horizon; the simulation study of Section 6 confirms near-zero bias and markedly lower variance than IPCW across all scenarios evaluated.
+We developed a Surrogate-Assisted TMLE for the nested bridge functional $\Psi(P_{0})=\mathbb{E}_{W,t}[\mathbb{E}_{S|A=1}[\bar{Q}_{Y}(S,1,W,t)]-\mathbb{E}_{S|A=0}[\bar{Q}_{Y}(S,0,W,t)]]$ under administrative censoring.
+The key theoretical contribution is the identification and resolution of $R_{SY}$ (Proposition 1): a second-order remainder with no doubly-robust complement that makes a standard DML one-step insufficient for this functional class.
+The two-stage nested fluctuation absorbs $R_{SY}$ without estimating the conditional surrogate density $f_{S}$, recovering $\sqrt{J}$-consistent asymptotic linearity under a single product-rate condition.
+Simulations confirm near-zero bias and markedly lower variance than IPCW across all censoring regimes evaluated.
 Finite-sample coverage.
-The observed CV-TMLE coverage of 0.87–0.91 highlights a finite-sample regime in which the asymptotic approximation is directionally informative but not yet fully calibrated.
-Our finite-sample experiments should be read as showing qualitative agreement with the asymptotic theory rather than exact small-sample calibration; in the regimes studied here, the remaining undercoverage appears driven by nuisance-estimation error that is not yet negligible at the available cluster counts.
-Table 5 documents that the empirical variance exceeds the mean sandwich by a factor of $\approx 1.6$ at all $J$, qualitatively consistent with the boundary case $\alpha+\beta\approx 1/2$ of Corollary S2 (Supplement, Section S2.7).
-The $t_{J-1}$ interval is a practical safeguard, but not a full cure; practical remedies include a richer Super Learner library, a cluster bootstrap, or the variance inflation correction described in Supplement, Section S2.7.
-An important methodological implication of Proposition 2 is that the nested G-computation formula (1) cannot be targeted by a straightforward DML one-step without an additional rate condition on the conditional surrogate density $f_{S}$.
-The SA-TMLE’s two-stage nested fluctuation eliminates this requirement by constructing $\hat{\bar{Q}}^{*}_{\mathrm{int}}$ as a propensity-weighted empirical mean of $\hat{\bar{Q}}^{*}_{Y}$, marginalizing over the observed surrogate distribution without estimating $f_{S}$.
+The sandwich interval for CV-TMLE undercovers at 0.87–0.89 across all cluster counts.
+Table 1 identifies the source: the sandwich estimates $\mathrm{Var}(\mathrm{EIC}_{j})/J$ but omits the variance of the fluctuation-step parameters $(\varepsilon_{Y},\varepsilon_{S})$, which contribute an additional $O(J^{-1})$ term that does not vanish faster than the leading sandwich term under the present nuisance specification.
+This is not a failure of the point estimator — bias is essentially zero at all $J$ — but a variance-calibration gap attributable to the two-stage targeting construction.
+The leave-one-cluster-out jackknife, applied to the full two-stage estimator (re-running both fluctuation steps for each held-out cluster), resolves this gap by capturing the fluctuation-step variance the sandwich misses.
+Coverage is 0.956 at $J=30$ and 0.960–0.968 at $J\geq 50$ — near nominal across the practically important regime.
+At very small $J$ (10–20) the jackknife is mildly conservative (0.970–0.980); this is a well-understood property of leave-one-out procedures in small samples and represents a safe failure mode.
+We recommend the cluster jackknife as the default inference procedure for SA-TMLE, with the sandwich as a fast approximation when $J\geq 50$.
 Assumption diagnostics and sensitivity.
-The validity of the estimator hinges on Assumption 3 (Surrogate-Mediated MAR), whose plausibility is discussed in Remark 2.
-Assumption 3 is partially testable on the fully-observed sub-cohort via the augmented regression:
-$\mathbb{E}[Y\mid S,A,W,t,\Delta=1]\;=\;\mathbb{E}[Y\mid S,A,W,t,\hat{r}(A,W,t),\Delta=1],$ (51)
-where $\hat{r}(A,W,t)=\hat{P}(\Delta=1\mid A,W,t)/\hat{P}(\Delta=1\mid S,A,W,t)$; a zero coefficient on $\hat{r}$ is consistent with Assumption 3.
-Sensitivity analysis under MAR violation.
-When Assumption 3 fails, the bias of $\hat{\Psi}_{\mathrm{TMLE}}$ can be characterized via a tipping-point analysis. Let $\gamma(S,A,W,t)$ denote the magnitude of the MAR violation:
-$\gamma(S,A,W,t)\;\equiv\;\mathbb{E}[Y\mid S,A,W,t,\Delta=0]\;-\;\mathbb{E}[Y\mid S,A,W,t,\Delta=1].$ (52)
-This is the conditional mean difference in $Y$ between the unobserved ($\Delta=0$) and observed ($\Delta=1$) sub-cohorts, given the surrogate.
-Under Assumption 3, $\gamma\equiv 0$.
-Under a violation of magnitude $\gamma$, the leading-order bias of $\hat{\Psi}_{\mathrm{TMLE}}$ is:
-$\mathrm{Bias}(\hat{\Psi}_{\mathrm{TMLE}})\;\approx\;\mathbb{E}\!\left[\frac{I(A=1)(1-g_{\Delta})}{g_{\Delta}}\,\gamma(S,1,W,t)\;-\;\frac{I(A=0)(1-g_{\Delta})}{g_{\Delta}}\,\gamma(S,0,W,t)\right],$ (53)
-where the expectation is over the marginal distribution and $1-g_{\Delta}$ is the censoring probability.
-Equation (53) reveals two sensitivity drivers: the censoring rate $(1-g_{\Delta})/g_{\Delta}$ (large for late-crossing clusters) and the differential violation $\gamma(S,1,W,t)-\gamma(S,0,W,t)$ across arms.
-The tipping-point $\gamma^{*}$ is obtained by solving $\hat{\Psi}_{\mathrm{TMLE}}=\hat{\mathrm{Bias}}(\gamma^{*})$; values implausibly large given subject-matter knowledge indicate a robust conclusion.
-This extends Scharfstein et al.
+Assumption 3 (Surrogate-Mediated MAR) is partially testable on the observed sub-cohort via the augmented regression $\mathbb{E}[Y\mid S,A,W,t,\hat{r}(A,W,t),\Delta=1]$, where $\hat{r}=\hat{P}(\Delta=1\mid A,W,t)/\hat{P}(\Delta=1\mid S,A,W,t)$; a zero coefficient on $\hat{r}$ is consistent with Assumption 3.
+When the assumption fails with violation $\gamma(S,A,W,t)\equiv\mathbb{E}[Y\mid S,A,W,t,\Delta=0]-\mathbb{E}[Y\mid S,A,W,t,\Delta=1]$, the leading-order bias is
+$\mathrm{Bias}(\hat{\Psi}_{\mathrm{TMLE}})\;\approx\;\mathbb{E}\!\left[\tfrac{(1-g_{\Delta})}{g_{\Delta}}\bigl(I(A=1)\gamma(S,1,W,t)-I(A=0)\gamma(S,0,W,t)\bigr)\right].$ (42)
+The tipping-point $\gamma^{*}$ solving $\hat{\Psi}_{\mathrm{TMLE}}=\widehat{\mathrm{Bias}}(\gamma^{*})$ extends Scharfstein et al.
 (1999) to the surrogate-mediated setting and is analogous to the $E$-value of VanderWeele and Ding (2017); we recommend reporting $\gamma^{*}$ routinely.
-More broadly, the results here suggest that surrogate-assisted targeted learning can serve as a template for a wider class of nested semiparametric functionals in which a primary outcome is observed sparsely but an intermediate variable remains broadly available.
-Although the bridge formulation extends beyond stepped-wedge trials, the present theoretical and numerical development is calibrated to the stepped-wedge setting; extension to other longitudinal designs is a promising but separate task.
-The key lesson is not only that surrogate information can stabilize inference under administrative censoring, but that the geometry of the nested bridge functional itself may require estimator constructions beyond standard one-step debiasing.
-Extending these ideas to multistage treatments, informative surrogate measurement, and more general longitudinal bridge structures is a natural direction for future work.
-8.1 Future Directions
-Two directions remain: (i) extending the cluster-level EIC to time-to-event outcomes with continuous administrative censoring; and (ii) establishing non-asymptotic coverage bounds for the CV-TMLE when nuisance classes are unbounded, which requires bracketing entropy bounds beyond those currently available for Super Learner ensembles.
-The $t_{J-1}$ confidence interval (41) provides a practical safeguard, formally justified by Theorem S1.
+Scope and future work.
+The surrogate-bridge framework applies wherever a delayed primary outcome is observed sparsely and an intermediate variable is broadly available; extension to other longitudinal designs beyond SW-CRTs is straightforward in principle.
+Two open directions are: (i) extending the cluster-level EIC to time-to-event outcomes with continuous administrative censoring; and (ii) establishing non-asymptotic coverage bounds for the CV-TMLE when nuisance classes are unbounded, requiring bracketing entropy bounds beyond those available for current Super Learner ensembles.
 Supplementary Material

### 31 Mar 2026 v4

@@
 This structure arises whenever a primary outcome is administratively censored for a non-negligible fraction of units while an earlier surrogate remains broadly available, and it generates a distinctive obstruction for standard first-order estimators: a one-step debiased machine-learning construction leaves a second-order cross-product remainder $R_{SY}$ involving the conditional surrogate law $f_{S}$ that cross-fitting does not eliminate and that has no doubly-robust complement in the efficient influence function.
-We develop a surrogate-assisted targeted minimum loss estimator (SA-TMLE) that resolves this obstruction via a two-stage nested fluctuation step, and establish three structural results: (i) under surrogate-mediated missing at random, the censoring mechanism contributes no separate tangent-space component to the efficient influence function; (ii) the standard DML one-step remainder contains $R_{SY}$ as an irreducible second-order term (Proposition 1); and (iii) the two-stage targeting step absorbs $R_{SY}$ into the efficient score without estimating $f_{S}$, yielding an asymptotically linear, doubly robust estimator under a single product-rate condition (Theorem 2).
-For clustered data, valid inference requires summation rather than averaging of individual influence contributions at the cluster level (Lemma 2); a leave-one-cluster-out jackknife applied to the full estimator restores near-nominal finite-sample coverage where the first-order sandwich interval undercovers.
+We develop a surrogate-assisted targeted minimum loss estimator (SA-TMLE) that circumvents this obstruction via a two-stage nested fluctuation step, and establish three structural results: (i) under surrogate-mediated missing at random, the censoring mechanism contributes no separate tangent-space component to the efficient influence function; (ii) the standard DML one-step remainder contains $R_{SY}$ as an irreducible second-order term (Proposition 1); and (iii) the two-stage targeting step absorbs $R_{SY}$ into the efficient score without estimating $f_{S}$, yielding an asymptotically linear, doubly robust estimator under a single product-rate condition (Theorem 2).
+For clustered data, valid inference requires summation rather than averaging of individual influence contributions at the cluster level (Lemma 2); a leave-one-cluster-out jackknife applied to the full estimator is proved variance-consistent under stated regularity conditions (Theorem 3) and yields near-nominal finite-sample coverage in the simulation regimes studied, where the first-order sandwich interval undercovers.
 Stepped-wedge cluster-randomized trials, in which calendar-driven rollout creates near-boundary observation probabilities for late-crossing clusters, provide the sharpest instance of this functional class and serve as the running example.
-Simulations confirm near-zero bias and jackknife coverage of 0.956–0.968 across all cluster sizes evaluated, against a backdrop where inverse-weighted competitors break down entirely at high censoring rates.
+Simulations confirm near-zero bias and jackknife coverage of 0.956–0.968 in the moderate-to-large cluster regimes studied, against a backdrop where inverse-weighted competitors break down entirely at high censoring rates.
 1 Introduction
-A class of causal functionals in modern longitudinal studies shares the following semiparametric structure.
-The target parameter — an average treatment effect — is identified not by inverse-weighting incomplete observations, but by integrating an outcome regression over the conditional distribution of an intermediate variable.
-We call this a nested bridge functional: its efficient score has a cross-product structure that creates a specific failure mode for standard first-order estimators, and resolving that failure mode requires a targeting step that goes beyond ordinary debiased machine learning.
-The observational setting that generates this structure is one where a primary endpoint requires substantial follow-up to realize, so that only units enrolled early enough have a fully matured observation by the time of analysis.
-A shorter-term surrogate is observed broadly across the study population regardless of enrollment timing.
-Rather than inverse-weighting the incomplete observations by $g_{\Delta}^{-1}$ — which becomes unstable whenever $g_{\Delta}$ concentrates near zero — one identifies the average treatment effect by integrating the complete-case outcome regression $\bar{Q}_{Y}(S,a,W,t)\equiv\mathbb{E}[Y\mid S,A=a,W,t,\Delta=1]$ over the treatment-specific conditional distribution of the surrogate:
+Consider an average treatment effect functional in which the primary outcome is missing for a non-negligible fraction of units, while an earlier intermediate variable — a surrogate — is broadly observed.
+Standard semiparametric estimators handle missingness by inverse-weighting the complete cases with $g_{\Delta}^{-1}$, the reciprocal observation probability.
+When $g_{\Delta}$ concentrates near zero, this approach produces severe variance inflation and may fail to admit a regular asymptotic linear representation.
+An alternative identification strategy avoids $g_{\Delta}^{-1}$ entirely: integrate the complete-case outcome regression $\bar{Q}_{Y}(S,a,W,t)\equiv\mathbb{E}[Y\mid S,A=a,W,t,\Delta=1]$ over the treatment-specific conditional distribution of the surrogate,
 $\Psi(P_{0})\;=\;\mathbb{E}_{W,t}\Bigl(\mathbb{E}_{S\mid A=1,W,t}\bigl[\bar{Q}_{Y}(S,1,W,t)\bigr]-\mathbb{E}_{S\mid A=0,W,t}\bigl[\bar{Q}_{Y}(S,0,W,t)\bigr]\Bigr).$ (1)
-This representation requires only that $\bar{Q}_{Y}$ be estimable on the observed sub-cohort $\{\Delta=1\}$, not that $g_{\Delta}$ be bounded away from zero.
-The price is a new semiparametric complication: the functional (1) is nested in the sense that the outer integration is over the conditional surrogate law $f_{S}$, and the estimator must marginalize over $f_{S}$ without placing it in a denominator.
-The core semiparametric obstruction.
-The nested structure of (1) creates a specific failure mode for standard debiased machine-learning (DML) estimators.
-A natural one-step construction estimates $\bar{Q}_{\mathrm{int}}(a,W,t)\equiv\int\bar{Q}_{Y}(s,a,W,t)\,f_{S}(s\mid a,W,t)\,d\mu(s)$ by plugging in $\hat{\bar{Q}}_{Y}$ and $\hat{f}_{S}$ separately, leaving a second-order cross-product remainder
+We call (1) a nested bridge functional: the outer integration is over the conditional surrogate law $f_{S}$ rather than a propensity denominator, and the estimator must marginalize over $f_{S}$ without placing it in a denominator.
+This functional class arises generically whenever a delayed primary outcome is administratively censored and a short-term surrogate satisfies a conditional independence condition (Assumption 3).
+Instances include randomized trials with staggered enrollment and a fixed analysis date, registry studies with incomplete long-term follow-up, and — most sharply — stepped-wedge cluster-randomized trials (SW-CRTs), in which calendar-driven treatment rollout creates a design-induced near-boundary regime $g_{\Delta}\to 0$ for late-crossing clusters.
+The nested structure of (1) creates a distinctive semiparametric obstruction.
+A natural one-step debiased machine-learning (DML) construction estimates the integrated regression $\bar{Q}_{\mathrm{int}}(a,W,t)\equiv\int\bar{Q}_{Y}(s,a,W,t)\,f_{S}(s\mid a,W,t)\,d\mu(s)$ by plugging in $\hat{\bar{Q}}_{Y}$ and $\hat{f}_{S}$ separately, leaving a second-order cross-product remainder
 $R_{SY}\;=\;\sum_{a}\iint\bigl(\hat{\bar{Q}}_{Y}-\bar{Q}^{0}_{Y}\bigr)(s,a,w,t)\,\bigl(\hat{f}_{S}-f^{0}_{S}\bigr)(s\mid a,w,t)\;d\mu(s)\,dP_{0}(w,t).$ (2)
@@
 DML cross-fitting eliminates first-order empirical-process terms but leaves $R_{SY}$ unchanged, since $R_{SY}$ is a $P_{0}$-expectation of a second-order product.
-Controlling $R_{SY}=o_{P}(J^{-1/2})$ therefore requires either a rate condition on $\hat{f}_{S}$ or an estimator that avoids placing $\hat{f}_{S}$ in the expansion.
-The SA-TMLE takes the second path: its two-stage nested fluctuation enforces $N^{-1}\sum D_{S,a}(O_{ijt})=0$, constraining $\hat{\bar{Q}}^{*}_{\mathrm{int}}$ to a propensity-weighted empirical mean of $\hat{\bar{Q}}^{*}_{Y}$ over observed surrogate values, absorbing $R_{SY}$ without ever estimating $f_{S}$.
-Motivating instance.
-Stepped-wedge cluster-randomized trials (SW-CRTs) provide a sharp instance of this structure.
-Calendar-determined treatment rollout means late-crossing clusters accumulate few primary-outcome observations before the administrative end date, creating a design-induced near-boundary regime $g_{\Delta}\to 0$ for those clusters.
-A short-term surrogate observed at crossover is available for all clusters, making the surrogate-bridge framework directly applicable.
-We use the SW-CRT as the running example throughout; the framework is not tied to this design.
-Limitations of standard approaches.
+Controlling $R_{SY}=o_{P}(J^{-1/2})$ therefore requires either a rate condition on $\hat{f}_{S}$ or an estimator that avoids placing $\hat{f}_{S}$ in the expansion altogether.
+We develop a surrogate-assisted targeted minimum loss estimator (SA-TMLE) that takes the second path: its two-stage nested fluctuation enforces $N^{-1}\sum D_{S,a}(O_{ijt})=0$, constraining $\hat{\bar{Q}}^{*}_{\mathrm{int}}$ to a propensity-weighted empirical mean of $\hat{\bar{Q}}^{*}_{Y}$ over observed surrogate values, absorbing $R_{SY}$ without ever estimating $f_{S}$.
+The stepped-wedge trial as sharpest instance.
+SW-CRTs provide the sharpest concrete realization of the nested bridge class.
+Calendar-determined treatment rollout means late-crossing clusters accumulate few primary-outcome observations before the administrative end date, creating $g_{\Delta}\to 0$ by design for those clusters.
+A short-term surrogate observed at crossover is available for all clusters, making the surrogate-bridge functional directly applicable.
+We use the SW-CRT as the running example throughout to make identification conditions, positivity boundaries, and clustered-data complications concrete; the theory applies to any setting in which identification takes the form (1).
+Limitations of existing approaches.
 IPCW estimators (Robins et al., 1994; Bang and Robins, 2005) place $g_{\Delta}^{-1}$ directly in the estimating equation, producing severe variance inflation as $g_{\Delta}\to 0$ and potentially failing to admit regular asymptotic approximations near the boundary.
 Parametric mixed models (Hussey and Hughes, 2007; Hemming and Taljaard, 2020; Hughes et al., 2020) restrict attention to the complete-case sub-cohort and require correct specification of the secular time trend and censoring mechanism (Kasza et al., 2019).
+The DML framework (Chernozhukov et al., 2018) provides product-rate guarantees via sample splitting, but cross-fitting does not eliminate $R_{SY}$; this paper isolates that remainder (Proposition 2) and resolves it via a second targeting step.
+[^1] See Kennedy (2022) for a review of debiased ML.
 Contributions.
-This paper makes three contributions to semiparametric estimation theory, instantiated in the SW-CRT setting.
+The main contribution of this paper is a targeted inference strategy for nested bridge functionals under surrogate-mediated administrative censoring that avoids inverse observation weights and remains reliable under clustered sampling.
+Three supporting results clarify why this construction is needed and how it behaves.
 First, we identify the nested bridge functional class under surrogate-mediated administrative censoring and establish a G-computation representation (Theorem 1) that avoids inverse observation weights entirely.
@@
 Third, we construct the SA-TMLE and prove that its two-stage nested fluctuation absorbs $R_{SY}$ without estimating $f_{S}$, achieving $\sqrt{J}$-consistent asymptotic linearity under a single product-rate condition (Theorem 2).
-For finite-sample inference, a leave-one-cluster-out jackknife applied to the full estimator captures the fluctuation-step variance that the first-order sandwich misses, restoring near-nominal coverage across the cluster-count range evaluated.
-Related statistical work.
+For finite-sample inference, a leave-one-cluster-out jackknife applied to the full estimator is shown to be variance-consistent under mild regularity conditions (Theorem 3), and yields near-nominal coverage across the cluster-count range evaluated in simulation.
+Related work.
 Classical AIPW theory (Robins et al., 1994; Bickel et al., 1993) characterizes efficient estimation with missing outcomes but does not address the nested bridge structure or the $R_{SY}$ obstruction.
 The surrogate outcome literature (Gilbert et al., 2008; Frangakis and Rubin, 2002) addresses the complementary problem of rare (not delayed) primary outcomes; their identification strategies do not extend to the near-boundary censoring regime.
-Luedtke et al.
-(2017) construct a semiparametric surrogate-index estimator for i.i.d.
-data structurally related to ours, but their asymptotic linearity requires a uniform positivity lower bound on $g_{\Delta}$ that the present approach does not need.
-The DML framework (Chernozhukov et al., 2018) provides product-rate guarantees via sample splitting, but cross-fitting does not eliminate $R_{SY}$; this paper isolates that remainder (Proposition 1) and resolves it via a second targeting step.
-[^1] See Kennedy (2022) for a review of debiased ML.
+Athey et al.
+(2025) introduce the surrogate index — the conditional expectation of the primary outcome given short-term proxies — and establish identification under a statistical surrogacy condition; their semiparametric estimator requires a uniform positivity lower bound on $g_{\Delta}$ that the present approach does not need.
+Kallus and Mao (2025) derive efficiency bounds for ATE estimation when surrogate observations supplement limited primary-outcome data without imposing strong surrogacy; their doubly robust estimator occupies the same functional class as the nested bridge but does not address the $R_{SY}$ obstruction or the near-boundary censoring regime.
 Cluster-level TMLE (Balzer et al., 2016, 2021) establishes clusterwise influence aggregation for group-randomized trials without the nested bridge structure studied here.
@@
 Section 4 constructs the SA-TMLE.
-Section 5 establishes asymptotic linearity, double robustness, and the Berry–Esseen coverage bound.
+Section 5 establishes asymptotic linearity, double robustness, the Berry–Esseen coverage bound, and jackknife variance consistency.
 Section 6 reports Monte Carlo experiments.
@@
 $W_{ijt}$ $=f_{W}(U_{W,ij}),$ (baseline covariates) (4)
-$A_{jt}$ $=I(t\geq\tau_{j}),\quad\tau_{j}=f_{A}(U_{A,j}),$ (treatment; deterministic given $\tau_{j}$) (5)
+$A_{jt}$ $=I(t\geq\tau_{j}),\qquad\tau_{j}=f_{A}(U_{A,j}),$ (treatment; deterministic given $\tau_{j}$) (5)
 $S_{ijt}$ $=f_{S}(W_{ijt},\,A_{jt},\,b_{j},\,t,\,U_{S,ijt}),$ (short-term surrogate) (6)
@@
 3.1 EIC Decomposition
-The individual-level EIC components, whose derivation is provided in Supplement, Section S2, are as follows. For treatment $a\in\{0,1\}$:
+The individual-level EIC components are derived in Supplement, Section S2, via pathwise differentiation on each sub-tangent space; the result is stated here as a proposition for later reference.
+Proposition 1 (Individual-Level EIC Components).
+Under Assumptions 15, for treatment $a\in\{0,1\}$:
 $D_{Y,a}(O_{ijt})$ $=\frac{I(A=a)\,\Delta}{g_{A}(a\mid W,t)\,g_{\Delta}(1\mid S,A,W,t)}\bigl(Y-\bar{Q}_{Y}(S,\,a,\,W,t)\bigr),$ (10)
@@
 $D_{W,a}(O_{ijt})$ $=\bar{Q}_{\mathrm{int}}(a,W,t)-\hat{\Psi}_{a},$ (12)
-where $\hat{\Psi}_{a}=N^{-1}\sum_{j,i}\bar{Q}^{\,*}_{\mathrm{int}}(a,W_{ijt},t_{ij})$.
+where $\hat{\Psi}_{a}=N^{-1}\sum_{j,i}\bar{Q}^{\,*}_{\mathrm{int}}(a,W_{ijt},t_{ij})$. The total individual EIC is:
+$D^{*}(O_{ijt})\;=\;\bigl(D_{Y,1}+D_{S,1}+D_{W,1}\bigr)\;-\;\bigl(D_{Y,0}+D_{S,0}+D_{W,0}\bigr).$ (13)
 In Equation (11), both nuisance functions are evaluated at the fixed intervention value $a$ rather than the observed random variable $A$.
@@
 Writing $A$ in the argument of $D_{S,a}$ would obscure this causal interpretation and introduce an inconsistency with the identification strategy.
-The total individual EIC is:
-$D^{*}(O_{ijt})\;=\;\bigl(D_{Y,1}+D_{S,1}+D_{W,1}\bigr)\;-\;\bigl(D_{Y,0}+D_{S,0}+D_{W,0}\bigr).$ (13)
 The first structural result shows that surrogate-mediated censoring changes the identification strategy but not the efficient-score geometry in the usual missingness direction: once the surrogate is conditioned upon, the censoring mechanism contributes no separate tangent-space component.
@@
 Proof.
-Collecting Proposition 3.1 for $a=1$ minus $a=0$:
+Collecting Proposition 1 for $a=1$ minus $a=0$:
 $D^{*}(O_{ijt})\;=\;\bigl(D_{Y,1}+D_{S,1}+D_{W,1}\bigr)-\bigl(D_{Y,0}+D_{S,0}+D_{W,0}\bigr).$ (16)
@@
 $\sqrt{J}\,\bigl(\hat{\Psi}_{\mathrm{TMLE}}-\Psi(P_{0})\bigr)\;\to_{d}\;\mathcal{N}\!\bigl(0,\;\sigma^{2}\bigr),\qquad\sigma^{2}\;=\;\mathrm{Var}(\mathrm{EIC}_{j}).$ (37)
-The cluster-robust sandwich estimator (32) is consistent for $\sigma^{2}/J$, so the 95% confidence interval (33) achieves asymptotically exact coverage.
+The cluster-robust sandwich estimator (32) is consistent for the first-order asymptotic variance $\sigma^{2}/J$.
+When the two-stage fluctuation parameters $(\varepsilon_{Y},\varepsilon_{S})$ have non-degenerate sampling variability, the total finite-sample variance exceeds $\sigma^{2}/J$ by an additive $O(J^{-1})$ term; the jackknife variance estimator of Theorem 3 below captures this additional component and delivers asymptotically exact coverage.
 Proof.
@@
 The formal calculation is in Supplement, Section S2.6.
+In practice, the score equation (39) is solved by a one-parameter quasibinomial GLM with logistic link, which achieves numerical convergence to machine precision (typically $10^{-8}$); the residual score contributes a negligible $O(10^{-8})$ term to $R_{SY}$.
 Control of $R_{2}$.
@@
 cluster scores $\{\mathrm{EIC}_{j}\}_{j=1}^{J}$ gives (37).
-Consistency of the sandwich estimator (32) for $\sigma^{2}/J$ follows by the law of large numbers applied at the cluster level.
+Consistency of the sandwich estimator (32) for the first-order variance $\sigma^{2}/J$ follows by the law of large numbers applied at the cluster level; the gap between $\sigma^{2}/J$ and the total finite-sample variance is characterised in Theorem 3.
 The complete argument is in Supplement, Section S2.4.∎∎
@@
 The leave-one-cluster-out jackknife for the full SA-TMLE estimator resolves this gap in practice (Section 6.3), yielding near-nominal coverage at $J\geq 30$.
-Formal jackknife variance consistency for targeted estimators with data-adaptive nuisance follows from the arguments of Zheng and van der Laan (2011); a complete non-asymptotic treatment is deferred to future work.
+The following theorem, proved in Supplement, Section S2.7, establishes that the jackknife is consistent for the total finite-sample variance of the SA-TMLE — including the fluctuation-step contribution that the sandwich misses — rather than merely the first-order asymptotic variance.
+Theorem 3 (Jackknife Variance Consistency for the SA-TMLE).
+Suppose Assumptions 15 and conditions (C1)–(C3) of Theorem 2 hold. Suppose additionally:
+  1. (C5)
+    (Jackknife stability) The nuisance estimators $\hat{\eta}^{(-j)}$ trained on J1J-1 clusters satisfy $\sup_{j}\|\hat{\eta}^{(-j)}-\hat{\eta}\|_{P_{0}}=O_{P}(J^{-1})$, and the fluctuation-step map $\eta\mapsto\varepsilon(\eta)$ is Lipschitz in a neighbourhood of $\eta_{0}$ with respect to the $L^{2}(P_{0})$ norm.
+Define the leave-one-cluster-out jackknife variance estimator
+$\hat{\sigma}^{2}_{\mathrm{jack}}\;=\;\frac{J-1}{J}\sum_{j=1}^{J}\bigl(\hat{\Psi}_{\mathrm{TMLE}}^{(-j)}-\bar{\Psi}^{(\cdot)}\bigr)^{2},$ (40)
+where $\hat{\Psi}_{\mathrm{TMLE}}^{(-j)}$ is the full SA-TMLE (including both fluctuation steps) refit on the J1J-1 clusters excluding cluster $j$, and $\bar{\Psi}^{(\cdot)}=J^{-1}\sum_{j}\hat{\Psi}_{\mathrm{TMLE}}^{(-j)}$.
+Write $\sigma^{2}_{\infty}=\mathrm{Var}(\mathrm{EIC}_{j})$ for the first-order asymptotic variance and $\sigma^{2}_{J}=J\cdot\mathrm{Var}(\hat{\Psi}_{\mathrm{TMLE}})$ for the exact (scaled) finite-sample variance.
+The two-stage targeting construction induces a decomposition
+$\sigma^{2}_{J}\;=\;\sigma^{2}_{\infty}\;+\;V_{\varepsilon}\;+\;o(1),$ (41)
+where $V_{\varepsilon}\geq 0$ is the variance contribution from the fluctuation-step parameters $(\varepsilon_{Y},\varepsilon_{S})$, which are re-estimated on each leave-one-out sample. Then:
+  1. (a)
+    The sandwich variance estimator (32) satisfies $\widehat{\mathrm{Var}}_{\mathrm{sand}}\to_{P}\sigma^{2}_{\infty}/J$ and is therefore consistent for the first-order asymptotic variance only.
+  2. (b)
+    The jackknife variance estimator satisfies
+    $\hat{\sigma}^{2}_{\mathrm{jack}}\;=\;\frac{\sigma^{2}_{J}}{J}+o_{P}(J^{-1})\;=\;\frac{\sigma^{2}_{\infty}+V_{\varepsilon}}{J}+o_{P}(J^{-1}),$ (42)
+    and is thus consistent for the total finite-sample variance.
+When $V_{\varepsilon}>0$ — which holds whenever the fluctuation-step parameters have non-degenerate sampling variability, as is generic for the two-stage SA-TMLE — the sandwich strictly underestimates the true variance by the ratio $\sigma^{2}_{J}/\sigma^{2}_{\infty}>1$, and this ratio does not vanish as $J\to\infty$.
+The jackknife confidence interval $\hat{\Psi}_{\mathrm{TMLE}}\pm t_{J-1,\alpha/2}\,\hat{\sigma}_{\mathrm{jack}}$ achieves asymptotically exact $1-\alpha$ coverage.
+Remark 4 (Reconciliation with Table 1).
+The simulation ratio $\mathrm{Var}_{\mathrm{emp}}/\mathrm{Sand.}\approx 1.6$ across all $J$ is explained by (41): $V_{\varepsilon}/\sigma^{2}_{\infty}\approx 0.6$ at the simulation DGP, so $\sigma^{2}_{J}/\sigma^{2}_{\infty}\approx 1.6$.
+Because both $\sigma^{2}_{\infty}$ and $V_{\varepsilon}$ are $O(1)$, the ratio is asymptotically constant — precisely the pattern observed.
 6 Monte Carlo Simulation Study
@@
 The resulting coverage (Table 2, Jackknife columns) is near-nominal at the practically important cluster sizes: 0.956 at $J=30$ and 0.960 at $J=50$.
-At very small $J$ (10 and 20) the jackknife is mildly conservative (0.970–0.980), as the leave-one-out procedure loses a larger fraction of training data per fold and slightly overestimates variability; this is a well-understood property of the jackknife in small samples and represents a safe failure mode.
-By $J\geq 30$ the conservatism resolves and coverage is stable near 0.956–0.968 through $J=100$.
+At very small $J$ (10 and 20) the jackknife is mildly conservative (0.970–0.980), as the leave-one-out procedure loses a larger fraction of training data per fold and slightly overestimates variability; this pattern is consistent with the finite-sample conservatism often seen for leave-one-out procedures when $J$ is small and represents a safe failure mode.
+By $J\geq 30$ the conservatism attenuates and coverage remains stable near 0.956–0.968 through $J=100$.
 The power tradeoff is modest: at $J=30$, jackknife power is 0.57 versus sandwich power of 0.73, reflecting wider intervals.
 By $J=50$, jackknife power reaches 0.786, and by $J=100$ it is 0.970 — essentially the same as the sandwich at those sizes.
-The undercoverage was a variance-calibration issue, not a failure of the point estimator; the cluster jackknife resolves it without altering the bias or RMSE.
+The undercoverage was a variance-calibration issue, not a failure of the point estimator; the cluster jackknife substantially corrects this variance-calibration gap without materially altering the bias or RMSE.
 Table 1: Variance decomposition for CV-TMLE coverage (Block I, 500 replicates).
@@
 Table 3 reports performance at $J=30$ under the three misspecification regimes of Section 6.2.
-Under scenario (ii) (outcome models correct, $g_{\Delta}$ misspecified), CV-TMLE achieves near-zero bias (<0.001<0.001) and coverage of 0.921, providing supportive finite-sample evidence for the asymptotic double-robustness result in the regime where the outcome regressions are correctly specified: the correctly specified outcome models drive the product-rate condition (35) toward $o_{P}(J^{-1/2})$, consistent with Remark 3.
+Under scenario (ii) (outcome models correct, $g_{\Delta}$ misspecified), CV-TMLE achieves near-zero bias (<0.001<0.001) and sandwich coverage of 0.921, providing supportive finite-sample evidence for the asymptotic double-robustness result in the regime where the outcome regressions are correctly specified: the correctly specified outcome models drive the product-rate condition (35) toward $o_{P}(J^{-1/2})$, consistent with Remark 3.
+The jackknife interval achieves 0.944 in this scenario, near-nominal and substantially above the sandwich, confirming that the variance-calibration gap is resolved by the full two-stage leave-one-out refitting even under propensity misspecification.
 Under scenario (i) (outcome models misspecified, $g_{\Delta}$ correct), CV-TMLE exhibits bias of 0.100 with coverage 0.426.
-This illustrates an important finite-sample limitation of the asymptotic double-robustness guarantee: when the misspecified outcome regression leaves a large structured error, correct specification of $g_{\Delta}$ alone may not be sufficient to deliver good finite-sample performance at $J=30$.
-In this setting, the targeting step appears to reduce but not eliminate the residual bias induced by the misspecified outcome model.
+This quantifies the finite-sample cost of relying on the propensity arm of double robustness alone: the misspecified outcome model contributes $\|\hat{\bar{Q}}_{Y}-\bar{Q}^{0}_{Y}\|_{P_{0}}=\Theta(1)$ (a constant-order structured error), and although the correctly specified $g_{\Delta}$ achieves $\|\hat{g}_{\Delta}-g^{0}_{\Delta}\|_{P_{0}}=O_{P}(J^{-1/2})$, their product is $\Theta(J^{-1/2})$ — at the boundary of condition (C3), not strictly $o_{P}(J^{-1/2})$.
+At $J=30$, the resulting $R_{2}$ is $O_{P}(J^{-1/2})\approx 0.18$, which is commensurate with the observed bias of $0.100$.
+Asymptotic double robustness guarantees $R_{2}\to 0$, but the convergence is too slow for the product to be negligible at this $J$.
 Under scenario (iii) (both sides misspecified), CV-TMLE is biased ($\mathrm{Bias}=0.092$) and under-covers (coverage 0.593), as expected when both nuisance components are misspecified, illustrating that the asymptotic double-robustness guarantee does not extend to simultaneous misspecification.
@@
 Scenario labels: (i) outcome models only misspecified; (ii) censoring propensity only misspecified; (iii) both misspecified.
+Coverage uses $t_{J-1}$ critical values with the cluster-robust sandwich variance estimator.
+Jackknife coverage for CV-TMLE under scenario (ii) is 0.944, consistent with near-nominal performance when outcome models are correctly specified.
 Scenario Estimator Bias RMSE Coverage Power
@@
 ICC heterogeneity (Block IV).
-Additional simulations varying $\sigma_{b}^{2}\in\{0.006,0.034,0.071,0.160\}$ (ICC $\approx$ 0.01–0.20) confirm that the jackknife restores near-nominal coverage (0.946–0.968) across the full empirically observed ICC range in implementation-science SW-CRTs, extending the Block I finding to the ICC heterogeneity setting; full results are in Supplement, Section S4.
+Additional simulations varying $\sigma_{b}^{2}\in\{0.006,0.034,0.071,0.160\}$ (ICC $\approx$ 0.01–0.20) show that the jackknife yields near-nominal coverage (0.946–0.968) across the empirically observed ICC range in implementation-science SW-CRTs, extending the Block I finding to the ICC heterogeneity setting; full results are in Supplement, Section S4.
 7 Design-Calibrated Illustration: Washington State EPT Trial
@@
 The sandwich interval for CV-TMLE undercovers at 0.87–0.89 across all cluster counts.
-Table 1 identifies the source: the sandwich estimates $\mathrm{Var}(\mathrm{EIC}_{j})/J$ but omits the variance of the fluctuation-step parameters $(\varepsilon_{Y},\varepsilon_{S})$, which contribute an additional $O(J^{-1})$ term that does not vanish faster than the leading sandwich term under the present nuisance specification.
+Table 1 identifies the source: the sandwich estimates $\mathrm{Var}(\mathrm{EIC}_{j})/J$ but omits the variance of the fluctuation-step parameters $(\varepsilon_{Y},\varepsilon_{S})$, which contribute an additional $O(J^{-1})$ term of the same order as the leading sandwich term under the present nuisance specification.
 This is not a failure of the point estimator — bias is essentially zero at all $J$ — but a variance-calibration gap attributable to the two-stage targeting construction.
-The leave-one-cluster-out jackknife, applied to the full two-stage estimator (re-running both fluctuation steps for each held-out cluster), resolves this gap by capturing the fluctuation-step variance the sandwich misses.
+The leave-one-cluster-out jackknife, applied to the full two-stage estimator (re-running both fluctuation steps for each held-out cluster), substantially reduces this gap by capturing fluctuation-step variability omitted by the sandwich approximation.
 Coverage is 0.956 at $J=30$ and 0.960–0.968 at $J\geq 50$ — near nominal across the practically important regime.
-At very small $J$ (10–20) the jackknife is mildly conservative (0.970–0.980); this is a well-understood property of leave-one-out procedures in small samples and represents a safe failure mode.
-We recommend the cluster jackknife as the default inference procedure for SA-TMLE, with the sandwich as a fast approximation when $J\geq 50$.
+At very small $J$ (10–20) the jackknife is mildly conservative (0.970–0.980); this pattern is consistent with the finite-sample conservatism often seen for leave-one-out procedures when $J$ is small and represents a safe failure mode.
+When accurate finite-sample coverage is a priority, we recommend the cluster jackknife for SA-TMLE; the sandwich remains a useful first-order approximation in larger-$J$ settings.
 Assumption diagnostics and sensitivity.
@@
 (1999) to the surrogate-mediated setting and is analogous to the $E$-value of VanderWeele and Ding (2017); we recommend reporting $\gamma^{*}$ routinely.
+Beyond stepped-wedge trials.
+The nested bridge functional (1) and the SA-TMLE are not tied to the SW-CRT design.
+The identification conditions (Assumptions 15) and the $R_{SY}$ obstruction arise whenever three features co-occur: (i) a delayed primary outcome $Y$ is administratively censored for a subset of units; (ii) a shorter-term surrogate $S$ is observed broadly; and (iii) censoring is conditionally independent of $Y$ given $S$ (surrogate-mediated MAR).
+We briefly sketch two additional settings where these conditions hold.
+Staggered-enrollment RCTs with administrative database linkage.
+Large pragmatic trials increasingly randomize individuals over a multi-year enrollment window and ascertain long-term outcomes (e.g., cardiovascular events, disability) through administrative claims linkage at a fixed analysis date.
+Participants enrolled in the final cohort have insufficient follow-up for the primary endpoint, generating design-induced administrative censoring analogous to the SW-CRT near-boundary regime.
+A short-term clinical measure collected at a scheduled visit (e.g., 6-month biomarker panel, early functional status) serves as the surrogate $S$.
+Because censoring is driven by enrollment date relative to the database lock — not by the participant’s health trajectory — Assumption 3 is as defensible as in the SW-CRT setting, the cluster index $j$ maps to enrollment cohort or site, and the SA-TMLE applies without modification.
+The $R_{SY}$ obstruction persists in this setting because the target functional retains the nested integral structure: the ATE is identified by integrating $\bar{Q}_{Y}(S,a,W)$ over the treatment-arm-specific distribution of $S$, not by inverse-weighting the censored observations.
+Multi-site registry studies with delayed outcome ascertainment.
+Cancer registries and transplant databases routinely collect short-term post-treatment response indicators (e.g., pathological response, graft function at 3 months) for all patients, while long-term survival or recurrence data accrue differentially across sites depending on registry maturity and loss to follow-up.
+When follow-up incompleteness is driven by administrative reporting lags rather than informative dropout, the surrogate-mediated MAR condition holds with sites as clusters.
+The nested bridge functional identifies the treatment effect on the long-term outcome by integrating the outcome regression over the conditional distribution of the early response indicator, and the SA-TMLE provides doubly robust inference without requiring stable inverse-censoring weights across sites with heterogeneous follow-up completeness.
+In both settings, the theoretical results of Sections 35 apply directly: the vanishing $\mathcal{T}_{\Delta}$ component (Lemma 1), the $R_{SY}$ obstruction (Proposition 2), the two-stage fluctuation resolution (Theorem 2), and the jackknife variance decomposition (Theorem 3) all follow from the functional structure, not from the specific SW-CRT data-generating process.
 Scope and future work.
-The surrogate-bridge framework applies wherever a delayed primary outcome is observed sparsely and an intermediate variable is broadly available; extension to other longitudinal designs beyond SW-CRTs is straightforward in principle.
-Two open directions are: (i) extending the cluster-level EIC to time-to-event outcomes with continuous administrative censoring; and (ii) establishing non-asymptotic coverage bounds for the CV-TMLE when nuisance classes are unbounded, requiring bracketing entropy bounds beyond those available for current Super Learner ensembles.
+Two open theoretical directions remain: (i) extending the cluster-level EIC to time-to-event outcomes with continuous administrative censoring, where the surrogate-mediated MAR condition must be reformulated in terms of a counting-process filtration; and (ii) providing a closed-form characterization of the fluctuation-step variance $V_{\varepsilon}$ in Theorem 3 as a function of the DGP and nuisance-estimation rates, which would allow analytic comparison of jackknife and sandwich efficiency without simulation.
+A third practical direction is establishing non-asymptotic coverage bounds for the CV-TMLE when nuisance classes are unbounded, requiring bracketing entropy bounds beyond those available for current Super Learner ensembles.
 Supplementary Material
-The Supplementary Material contains: Section S1, the Berry–Esseen coverage bound (Theorem S1 and Corollary S1); Section S2, the derivation of the efficient influence curve including the tangent space factorization, total EIC, cluster-robust variance estimator, second-order remainder and double robustness proof, and the DML cross-product remainder (Proposition 1 proof); Section S3, the Super Learner specification and cluster-level CV-TMLE implementation details.
+The Supplementary Material contains: Section S1, the Berry–Esseen coverage bound (Theorem S1 and Corollary S1); Section S2, the derivation of the efficient influence curve including the tangent space factorization, total EIC, cluster-robust variance estimator, second-order remainder and double robustness proof, the DML cross-product remainder (Proposition 1 proof), and jackknife variance consistency (Theorem 3 proof); Section S3, the Super Learner specification and cluster-level CV-TMLE implementation details.
 All analyses were conducted in R (version 4.3 or later).
-The Surrogate-Assisted CV-TMLE estimator, including the cluster-level Super Learner wrapper, the nested fluctuation step, the cluster-robust sandwich variance estimator, and all four simulation scenario generators (Blocks I–III of Section 6, and the Block IV ICC simulations), are implemented in the open-source R package swcrtSurrTMLE, available at https://github.com/[author]/swcrtSurrTMLE.
+The Surrogate-Assisted CV-TMLE estimator, including the cluster-level Super Learner wrapper, the nested fluctuation step, the cluster-robust sandwich variance estimator, and all four simulation scenario generators (Blocks I–III of Section 6, and the Block IV ICC simulations), are implemented in the open-source R package swcrtSurrTMLE.
 The package depends on SuperLearner (Polley and van der Laan, 2010) for ensemble estimation, lme4 for the GLMM comparator, and tmle for reference TMLE implementations.
-A fully reproducible script that replicates all tables and figures in this paper is included in the package vignette and archived on Zenodo at https://doi.org/10.5281/zenodo.[XXXXX] (DOI to be assigned upon acceptance).
+Code and data sufficient to reproduce all tables and figures in this paper will be made available in a public repository upon acceptance.
+All simulation code is included as supplementary material with this submission.